### v1

# Improved Approximation Algorithms for Chromatic and Pseudometric-Weighted Correlation Clustering

Abstract
Correlation Clustering (CC) is a foundational problem in unsupervised learning that models binary similarity relations using labeled graphs. While classical CC has been widely studied, many real-world applications involve more nuanced relationships, either multi-class categorical interactions or varying confidence levels in edge labels. To address these, two natural generalizations have been proposed: Chromatic Correlation Clustering (CCC), which assigns semantic colors to edge labels, and pseudometric-weighted CC, which allows edge weights satisfying the triangle inequality. In this paper, we develop improved approximation algorithms for both settings. Our approach leverages LP-based pivoting techniques combined with problem-specific rounding functions. For the pseudometric-weighted correlation clustering problem, we present a tight $\frac{10}{3}$-approximation algorithm, matching the best possible bound achievable within the framework of standard LP relaxation combined with specialized rounding. For the Chromatic Correlation Clustering (CCC) problem, we improve the approximation ratio from the previous best of $2.5$ to $2.15$, and we establish a lower bound of $2.11$ within the same analytical framework, highlighting the near-optimality of our result.
1 Introduction
Clustering is a fundamental task in unsupervised learning, where the goal is to partition a set of objects into groups based on their pairwise relationships. One prominent problem in this domain is Correlation Clustering (CC) bansal2004correlation , which models binary similarity/dissimilarity between items using an edge-labeled graph: similar pairs are marked with a ‘$+$’ label and dissimilar pairs with a ‘--’. The objective is to partition the nodes to minimize disagreements—i.e., cases where the partitioning contradicts the edge labels. Due to its flexibility in not requiring a predefined number of clusters, CC has been widely utilized in various areas such as detecting communities in networks Chen2012 , inferring labels from user interactions Agrawal2009 ; Chakrabarti2008 and resolving ambiguous entities Kalashnikov2008 .
However, classic CC models only binary relationships, which is insufficient for many practical applications. For example, in a social network, edges may represent diverse relationship types such as “colleague,” “classmate,” or “family.” To address this limitation, Bonchi et al. bonchi2012chromatic introduced the Chromatic Correlation Clustering (CCC) problem, which generalizes CC to multi-class categorical settings. In CCC, the input is an edge-colored graph where each color represents a different relationship type. The goal is to cluster the nodes and assign a single color to each cluster such that the number of disagreements—edges whose color does not match the cluster’s assigned color, or edges that should be separated—is minimized. CCC has wide applications in link classification, entity resolution, and clustering in bioinformatics bonchi2012chromatic ; anava2015chromatic ; klodt2021color .
In parallel, another important generalization of CC is the weighted correlation clustering problem, where edges are associated with weights reflecting the reliability or cost of violating a given label. When weights are unrestricted, obtaining a constant-factor approximation is known to be hard (under the Unique Games Conjecture) khot2002ugc . However, when edge weights form a pseudometric—i.e., they satisfy the triangle inequality—constant-factor approximations become feasible. This weighted setting more faithfully models scenarios where not all edges are equally trustworthy.
1.1 Related Works
The Correlation Clustering (CC) problem has been widely studied since its introduction ben-dor1999clustering , and it is known to be APX-hard, leading to efforts to develop approximation algorithms. Early work by Bansal, Blum, and Chawla introduced a constant-factor approximation algorithm bansal2004correlation . Charikar et al. CHARIKAR2005360 improved this to a 4-approximation using linear programming. Ailon, Charikar, and Newman then introduced the Pivot algorithm ailon2008aggregating , which achieved a 3-approximation in linear time. Chawla et al. chawla2015near further improved this to 2.06 using more refined LP-rounding techniques. More recently, researchers have surpassed the 2-approximation barrier. Cohen-Addad, Lee, and Newman cohenaddad2022sherali used the Sherali-Adams hierarchy to develop a $(1.994+\varepsilon)$-approximation, while Cohen-Addad et al. cohenaddad2023preclustering proposed preclustering, which improved the approximation to $(1.73+\varepsilon)$. The most recent breakthrough by Cao et al. cao2024clusterlp introduced the cluster LP, which unifies all known LP relaxations for CC. They show that this can be approximated efficiently using preclustering, achieving a $(1.437+\varepsilon)$-approximation, the best known guarantee for CC so far. In a more recent work cao2024clusterlp , they introduced a new approach to find a feasible solution for the cluster LP in sublinear time.
Chromatic Correlation Clustering (CCC) is an extension of the classical Correlation Clustering (CC) problem, where edge colors represent different types of relationships. Bonchi et al. bonchi2012chromatic introduced CCC and proposed a heuristic algorithm without approximation guarantees. Anava et al. anava2015chromatic presented a 4-approximation algorithm based on LP rounding and also proposed two practical methods: Reduce and Cluster (RC), with an approximation ratio of 11, and a heuristic named Deep Cluster (DC). Klodt et al. klodt2021color showed that the Pivot algorithm by Ailon et al. ailon2008aggregating , originally designed for standard CC, provides a 3-approximation guarantee for CCC. They also established that the RC algorithm achieves a 5-approximation for CCC. However, Pivot’s color-agnostic nature limits its performance in practice. To address this, Klodt et al. proposed a heuristic called Greedy Expansion (GE), which performs well empirically despite lacking theoretical guarantees. More recently, Xiu et al. XiuHTCH22 developed a 2.5-approximation algorithm for CCC based on a linear programming approach, improving upon the previous best-known ratio. They also introduced a greedy heuristic that achieves strong empirical results.
In modern data analysis, correlation clustering must often be performed under computational constraints such as limited memory or streaming access to data. Consequently, substantial research has focused on crafting clustering algorithms specifically tailored for dynamic, streaming, online, and distributed settings (lattanzi2017consistent, ; fichtenberger2021, ; jaghargh2019, ; cohen2019fully, ; guo2021distributed, ; cohen2022online, ; assadi2022sublinear, ; lattanzi2021parallel, ; behnezhad2022, ; behnezhad2023, ; bateni2023, ; cohen2024dynamic, ; braverman2025fully, ).
1.2 Our Results
Our Contributions.
In this work, we present improved approximation algorithms for both the CCC and pseudometric-weighted CC problems.
  • For the pseudometric-weighted correlation clustering problem, we develop a refined LP-based pivoting algorithm that achieves a tight $\frac{10}{3}$-approximation. We further prove that this approximation factor is optimal within the standard LP relaxation framework combined advanced rounding functions.
  • For the Chromatic Correlation Clustering (CCC) problem, we enhance the LP-based method through a new analysis that yields a $2.15$-approximation, improving upon the previous best bound of $2.5$ by Xiu et al. XiuHTCH22 . We also establish a lower bound of $2.11$ within the same analytical framework, underscoring the near-optimality of our approach.
Both results are obtained by extending and unifying the triple-based analysis of LP-rounding schemes. Our work improves the theoretical guarantees for two natural and practically motivated generalizations of correlation clustering and contributes new insights into their structural and algorithmic properties.
Technical Overview.
Our algorithms for both Chromatic Correlation Clustering (CCC) and pseudometric-weighted Correlation Clustering (CC) build on linear programming (LP) relaxations and a unified triple-based rounding framework chawla2015near . Below, we outline the key technical insights:
Pseudometric-Weighted CC: The upper bound for the approximation factor $10/3$ is derived using the LP-based Pivot algorithm and a more careful rounding function. For the lower bound, By assuming the existence of an $\alpha$-approximation and analyzing carefully constructed hard instances, the technique derives necessary conditions that any rounding function must satisfy. These conditions expose inherent conflicts, demonstrating that $\alpha$ cannot be arbitrarily small. In particular, the analysis establishes that $\alpha$ must be at least $\frac{10}{3}$. The core idea is to identify instance configurations that induce contradiction between the properties required of the rounding functions, ultimately leading to this lower bound on $\alpha$.
Chromatic Correlation Clustering (CCC): Building on the LP formulation introduced by Xiu et al. XiuHTCH22 , which jointly encodes fractional cluster membership and color assignments. The decoupling of color assignment from cluster formation, allowing us to preserve color structure without entangling it with clustering decisions. Using a triple-based analysis, we introduce tailored rounding functions—particularly for neutral edges—to better align the rounding behavior with the LP’s structure and avoid overcounting. This careful handling of intra-color, conflicting, and neutral edges reduces the approximation factor from 2.5 to 2.15. Our lower bound analysis builds on the general triple-based framework, augmented with structural insights specific to the LP-CCC algorithm and its associated LP solution. We carefully define the cost and LP contribution of each edge type—particularly neutral edges—and construct adversarial instances that expose limitations of any rounding strategy.
Paper Organization.
The remainder of the paper is structured as follows: Section 2 introduces the problem formulations and LP relaxations for both pseudometric-weighted and chromatic correlation clustering. Section 3 presents our approximation algorithms and outlines their design. Section 4 defines the rounding functions used in the LP-based algorithms. Section 5 provides a detailed triple-based analysis of the approximation guarantees. We conclude with a summary and discussion in Section 6.
2 Preliminaries
The correlation clustering (CC) problem takes as input a signed undirected graph $G=(V,E=E^{+}\uplus E^{-})$, where each edge $e=uv\in E$ is assigned a sign ‘$+$’ or ‘--’, described by $e\in E^{+}$ or $e\in E^{-}$. The objective is to find a partition of the nodes such that the number of disagreements—i.e., negative edges within the same cluster and positive edges between different clusters—is minimized. In other words, the cost of the clustering $\mathcal{C}$ is as follows:
$\text{obj}(\mathcal{C}):=\sum_{uv\in E^{+}}x_{uv}+\sum_{uv\in E^{-}}(1-x_{uv}),$
where $x_{uv}=0$ indicates that there exists $C\in\mathcal{C}$ such that $u,v\in C$, and $x_{uv}=1$ otherwise.
CC has a standard LP relaxation leveraging the viewpoint on $x$ as a discrete metric between partitions. Since the $x$ above satisfies the triangle inequality, we can relax the range of $x$ from $\{0,1\}$ to $[0,1]$, resulting in the following LP:
minimize $\sum_{uv\in E^{+}}x_{uv}+\sum_{uv\in E^{-}}(1-x_{uv})$ (CC-LP)
subject to $x_{uv}+x_{vw}\geq x_{wu},$ (1)
$x_{uv}\in[0,1].$ (2)
The integrality gap of CC-LP on a complete graph is known to be $2$ CHARIKAR2005360 , which indicates that the standard LP-based algorithm cannot obtain a better approximation factor below $2$.
2.1 Pseudometric-weighted Correlation Clustering
The weighted Correlation Clustering (CC) problem is a generalization of the classical CC problem in which each edge is associated with a nonnegative violation cost. Specifically, for each edge $uv$, a weight $w_{uv}\geq 0$ is provided, and violating the edge’s label (either ‘$+$’ or ‘--’) incurs a penalty of $w_{uv}$. This differs from the standard setting, where all violations incur a uniform cost of $1$. The weighted variant allows us to encode edge-wise reliability: when $w_{uv}$ is large, it is more reasonable to follow the given label between $u$ and $v$.
However, assuming the Unique Games Conjecture, no $O(1)$-approximation algorithm exists for the general weighted case DEMAINE2006172 . An exception occurs when the weight function satisfies the triangle inequality, i.e., the weights form a pseudometric. In this pseudometric-weighted setting, a constant-factor approximation is known DBLP:conf/soda/CharikarG24 . Following the analysis of Charikar and Gao with $L=2$ yields an approximation factor of $\overline{B}_{\text{HR}}+\frac{1}{\frac{1}{3}}\leq\frac{4}{\frac{1}{3}}+2(L-1)+\frac{1}{\frac{1}{3}}=17,$ since the second type of charge occurs at most $L-1=1$ time in the charging scheme. The following is a natural LP relaxation of the weighted CC problem, extending (CC-LP):
minimize $\sum_{uv\in E^{+}}w_{uv}\cdot x_{uv}+\sum_{uv\in E^{-}}w_{uv}\cdot(1-x_{uv})$ (wCC-LP)
subject to $x_{uv}+x_{vw}\geq x_{wu},$ (3)
$x_{uv}\in[0,1].$ (4)
Here, the variable $x$ can be viewed as defining a pseudometric over the vertex set, representing the distance between clusters. Since CC on bipartite graphs has an integrality gap of 3 chawla2015near , and is a special case of pseudometric-weighted CC, the LP relaxation (wCC-LP) for pseudometric-weighted CC also has an integrality gap of at least 3.
2.2 Chromatic Correlation Clustering Problem
The Chromatic Correlation Clustering (CCC) problem is a variant of the classical CC problem in which each cluster is additionally assigned a color bonchi2012chromatic . The input includes a set of $L$ possible colors, as well as a special color $\gamma$ that denotes that two vertices should not be placed in the same cluster—analogous to a negative (‘--’) edge in the classical CC setting. When $L=1$ (i.e., a single cluster color), CCC reduces to the standard CC problem.
The following is a linear programming (LP) relaxation of the CCC problem bonchi2012chromatic :
minimize $\sum_{\phi(uv)\neq\gamma}x_{uv}^{\phi(uv)}+\sum_{\phi(uv)=\gamma}\sum_{c\in L}(1-x_{uv}^{c})$ (CCC-LP)
subject to $x_{uv}^{c}\geq x_{u}^{c},\,x_{v}^{c},$ (5)
$x_{uv}^{c}+x_{vw}^{c}\geq x_{wu}^{c},$ (6)
$\sum_{c\in L}x_{u}^{c}=|L|-1,$ (7)
$x_{u}^{c},\,x_{uv}^{c}\in[0,1].$ (8)
Here, the variables $x_{u}^{c}$ and $x_{uv}^{c}$ are soft assignments:
  • $1-x_{u}^{c}\in[0,1]$ represents the fractional assignment of vertex $u$ to a cluster of color $c$.
  • $1-x_{uv}^{c}\in[0,1]$ indicates the fractional agreement between vertices $u$ and $v$ under color $c$.
These variables measure the likelihood of vertices or edges being assigned to a color, with $\{1-x_{u}^{c}\}_{c\in L}$ forming a probability distribution over the colors assigned to vertex $u$, subject to constraints (7) and (8).
There is also a geometric interpretation of these variables. Consider $L$ discrete pseudometric spaces $(V_{c},d_{c})$ where $V_{c}=\{u_{c}:u\in V\}$, and vertex $u$ is connected to $u_{c}$ with a link of length $x_{u}^{c}$. Then, $x_{uv}^{c}$ represents the bottleneck distance between $u$ and $v$, conditioned on traversing the auxiliary connections $u\rightarrow u_{c}$ and $v\rightarrow v_{c}$. This view generalizes the classical CC setting, where the cluster-wise discrete metric can be regarded as a special case of bottleneck distances.
Since CCC generalizes the standard CC problem, the integrality gap of (CCC-LP) is at least as large as that of (CC-LP), which is 2.
3 Approximation Algorithm
Building on the LP formulations introduced in the previous sections, we now present approximation algorithms for both pseudometric-weighted CC and CCC settings.
Algorithm 1 LP-Pivot
Input: Graph $G=(V,\,E=E^{+}\uplus E^{-}\uplus E^{\circ})$, LP solution $\{x_{uv}\}_{uv\in E}$.
Output: Clustering $\mathcal{C}$ of $V$.
Pick a pivot $v\in V$ uniformly at random.
Set $C=\{v\}$.
for $u\in V\backslash\{v\}$do
     Set $p_{uv}$ as following:
$p_{uv}=\begin{cases}f^{+}(x_{uv}),&uv\in E^{+};\\ f^{-}(x_{uv}),&uv\in E^{-};\\ f^{\circ}(x_{uv}),&uv\in E^{\circ}.\end{cases}$
     Update $C\leftarrow C\cup\{u\}$ with probability $1-p_{uv}$.
end for
return $\mathcal{C}=\{C\}\cup\textsc{LP-Pivot}(G|_{V\backslash C},x|_{V\backslash C})$.
The algorithm for the pseudometric-weighted CC problem extends the classical LP-based pivoting method. It uses the LP-Pivot algorithm, given in Algorithm 1, which recursively selects a pivot vertex and forms clusters via randomized rounding based on LP distances, along with the specific rounding functions (11). For each edge $uv$ incident to the pivot $v$, the probability of $u$ being merged with $v$ into a same cluster is determined by not only on the LP solution $x_{uv}$, but also on the edge’s type as specified in the input. Rounding functions translate the LP solution ${x_{uv}}$ to a rounding probability to guarantee a constant-factor approximation by controlling the expected clustering cost.
The LP-CCC algorithm for CCC, given in Algorithm 2, first partitions the vertices according to their LP-derived color distributions. It then applies the LP-Pivot algorithm, with edge types partitioned by color. CCC-specific rounding functions (equations 12 and 13) are used to account for the color structure, and the final clustering is obtained by combining the color class clusters.
Algorithm 2 LP-CCC
Input: Graph $G=(V,\,E)$, color function $\phi:E\rightarrow L\cup\{\gamma\}$, LP solution $\{x_{u}^{c}\}_{u\in V,\,c\in L}$ and $\{x_{uv}^{c}\}_{uv\in E,\,c\in L}$.
Output: Clustering $\mathcal{C}$ of $V$, Coloring function $\Phi:\mathcal{C}\rightarrow L$.
Initialize $\mathcal{C}=\emptyset,\,S_{c}=\emptyset$ for all $c\in L$.
for $u\in V$ do
     if $\exists c\in L$ s.t. $x_{u}^{c}<\frac{1}{2}$then
         Update $S_{c}\leftarrow S_{c}\cup\{u\}$.
     else
         Update $\mathcal{C}\leftarrow\mathcal{C}\cup\{\{u\}\}$.
         Assign $\Phi(\{u\})$ as an arbitrary color.
     end if
end for
for $c\in L$ do
     $G_{c}=(S_{c},\,E_{c}=E_{c}^{+}\uplus E_{c}^{-}\uplus E_{c}^{\circ})$, where $E_{c}=E\cap\binom{S_{c}}{2}$,
     and $E^{+}\uplus E^{-}\uplus E^{\circ}$ is defined as a partition by color $c,\,\gamma,L\backslash\{c\}$ respectively.
     Set $\mathcal{C}_{c}=\textsc{LP-Pivot}(G_{c},x^{c}|_{E_{c}})$.
     Update $\mathcal{C}\leftarrow\mathcal{C}\cup\mathcal{C}_{c}$.
     Assign $\Phi(C)=c$ for all $C\in\mathcal{C}_{c}$.
end for
return $\mathcal{C},\,\Phi$.
4 Rounding Functions
The effectiveness of the LP-Pivot and LP-CCC algorithms critically depends on the choice of rounding functions used in the clustering process. Rounding functions $f^{+},\,f^{-},\,f^{\circ}:[0,1]\rightarrow[0,1]$ convert the LP value $x_{uv}$ to the non-selection probability $p_{uv}$ chawla2015near . The sign of the edge $uv$—either ‘$+$’, ‘--’, or $\circ$’—determines which rounding function is applied. The sign $\circ$’ indicates that the edge does not belong to $E$. The following natural conditions are imposed on any rounding function $f$:
$f(0)=0,\,f(1)=1;$ (9)
$x<y\Rightarrow f(x)\leq f(y).$ (10)
Condition 9 is not only intuitive but also necessary in certain cases, such as ensuring $f^{+}(0)=0$ and $f^{-}(1)=1$. Other constraints are not required in the proofs of Theorems 1 and 3, which thus provide lower bounds on the approximation factors for ‘general’ rounding functions.
Lemma 1.
The LP-Pivot algorithm achieves a constant-factor approximation in expectation only if $f^{+}(0)=0$ and $f^{-}(1)=1$.
Proof.
We prove it by contradiction on some graph instances.
Case 1 $f^{+}(0)=0.$ Consider $G=(V,E=\binom{V}{2}=E\uplus\emptyset\uplus\emptyset)$. The optimal clustering is $\mathcal{C}^{*}=\{V\}$, satisfying $\text{obj}(\mathcal{C}^{*})=0$, and the optimal LP solution is $x^{*}\equiv 0$.
Suppose $f^{+}(0)>0$. Then $\Pr[\textsc{LP-Pivot}(G,0)\neq\mathcal{C}^{*}]>0$. Since $\text{obj}(\mathcal{C})>0$ if and only if $\mathcal{C}\neq\mathcal{C}^{*}$, this leads to a contradiction with the assumption of the expected constant factor approximation.
Case 2 $f^{-}(1)=1.$ Consider $G=(V,E=\binom{V}{2}=\emptyset\uplus E\uplus\emptyset)$. The optimal clustering is $\mathcal{C}^{*}=\{\{v\}:v\in V\}$. The following arguments are similar to Case 1. ∎
Different variants of the CC problem may use different rounding functions. In this paper, we provide rounding functions for both the pseudometric-weighted CC and CCC problems.
4.1 Pseudometric-weighted Correlation Clustering
We propose the following rounding functions that yield a tight approximation factor:
$f^{+}(x)=f^{-}(x)=\begin{cases}0,&x<0.4;\\ \frac{5}{3}x,&0.4\leq x<0.6;\\ 1,&x\geq 0.6.\end{cases}$ (11)
With these functions, the algorithm achieves an expected approximation factor of $10/3$. Moreover, no other rounding function can improve this factor, as shown in Section 5.1.
4.2 Chromatic Correlation Clustering
We further consider the rounding functions $f^{+},\,f^{-}$ from Chawla et al. chawla2015near , which yield a $2.06$-approximation for classical CC, and introduce a new function $f^{\circ}$ to handle $\circ$-edges for CCC:
$f^{+}(x)=\begin{cases}0,&x<0.19;\\ \left(\frac{x-0.19}{0.5095-0.19}\right)^{2},&0.19\leq x<0.5095;\\ 1,&x\geq 0.5095,\end{cases}\quad f^{-}(x)=x,$ (12)
$f^{\circ}(x)=\begin{cases}1.7x,&x<0.5;\\ 0.3x+0.7,&x\geq 0.5.\end{cases}$ (13)
This rather intricate function was computed to not intersect with analytic bounds that violate an approximation factor of $\alpha=2.15$, as illustrated in Figure 1.
Figure 1: The region where $f^{\circ}$ violates the $\alpha=2.15$-approximation for CCC using the proposed $f^{\circ}$ defined in (13).
4.3 Comparison with Prior Work
Pseudometric-weighted CC: The LP-UMVD-Pivot algorithm, recently proposed by Charikar and Gao DBLP:conf/soda/CharikarG24 for the Ultrametric Violation Distance (UMVD) problem, follows a pivoting-based rounding strategy applied to an LP relaxation. When the number of distinct pairwise distances between elements, denoted by $L$, is equal to 2, their algorithm can be viewed as a special case of our LP-Pivot framework. In this setting, the two distances $d_{1}$ and $d_{2}$ correspond to the ‘--’ and ‘$+$’ labels, respectively, used in our rounding procedure. The rounding functions used are:
[…]
Here, $\alpha$ and $\beta$ are fixed algorithmic parameters. For the pseudometric-weighted CC problem, they are set to $\alpha=\frac{1}{3}$ and $\beta=0$, and $f^{\circ}$ is unused.
With triple-based analysis, this choice yields an approximation factor of $6$, as shown in the Appendix.
Chromatic CC: The LP-based pivoting algorithm by Xiu et al. XiuHTCH22 uses LP values directly as probabilities, corresponding to the following rounding functions:
$f^{+}(x)=f^{-}(x)=f^{\circ}(x)=x.$
This setting is known to achieve an approximation factor of $2.5$.
5 Triple-based Analysis
To complete the analysis of the algorithm, it suffices to show that for every triple of vertices $u,v,w\in V$, the expected cost incurred by the algorithm, denoted $ALG(uvw)$, is at most a factor $\alpha$ times the corresponding LP cost $LP(uvw)$. That is,
$ALG(uvw)\leq\alpha\cdot LP(uvw).$
If the inequality holds for every triple, then the total expected cost of the algorithm is at most $\alpha\cdot LP$. To show this, the analysis expresses the expected algorithmic cost and LP cost as averages over all possible pivot choices and vertex triples. Specifically, it defines:
  • $e.cost_{w}(u,v)$: the expected cost of violating constraint $(u,v)$, conditioned on pivot $w$,
  • $e.lp_{w}(u,v)$: the expected LP charge of edge $(u,v)$, conditioned on pivot $w$.
5.1 Pseudometric-weighted Correlation Clustering
In the CC setting, we use the function $\mathcal{C}$, as defined in chawla2015near , to measure the gap:
$\mathcal{C}(x_{uv},x_{vw},x_{wu},p_{uv},p_{vw},p_{wu})=\alpha\cdot LP(uvw)-ALG(uvw),$
where
$ALG(uvw)$ $=e.cost_{w}(uv)+e.cost_{u}(vw)+e.cost_{v}(wu),$
$LP(uvw)$ $=e.lp_{w}(uv)+e.lp_{u}(vw)+e.lp_{v}(wu),$
and
$e.cost_{w}(u,v)$ =$=\begin{cases}p_{uw}(1-p_{vw})+(1-p_{uw})p_{vw},&uv\in E^{+};\\ (1-p_{uw})(1-p_{vw}),&uv\in E^{-},\end{cases}$
$e.lp_{w}(u,v)$ =$=\begin{cases}(1-p_{uw}p_{vw})x_{uv},&uv\in E^{+};\\ (1-p_{uw}p_{vw})(1-x_{uv}),&uv\in E^{-}.\end{cases}$
In the weighted CC setting, edge weights further influence the value of $\mathcal{C}$:
$\mathcal{C}(x_{uv},x_{vw},x_{wu},p_{uv},p_{vw},p_{wu},w_{uv},w_{vw},w_{wu})=\alpha\cdot LP(uvw)-ALG(uvw),$
with the definition for $e.cost$ and $e.lp$ remains the same; the classical CC corresponds to $(w_{uv},w_{vw},w_{wu})=(1,1,1)$.
Under the pseudometric constraint on weights $w$, we can reduce the number of cases to consider in the analysis.
Lemma 2.
If $\alpha\cdot LP(uvw)-ALG(uvw)\geq 0$ holds for weight configurations $(w_{uv},w_{vw},w_{wu})\in\{(1,1,0),\,(1,0,1),\,(0,1,1)\}$, then the inequality also holds for any configuration $(w_{uv},w_{vw},w_{wu})$ satisfying the triangle inequality.
Proof.
Let all $x_{uv},\,x_{vw},\,x_{wu},\,p_{uv},\,p_{vw},\,p_{wu}$ be fixed. $ALG(uvw)$ and $LP(uvw)$ can be written as
$ALG(uvw)=w_{uv}\cdot e.cost_{w}(uv)+w_{vw}\cdot e.cost_{u}(vw)+w_{wu}\cdot e.cost_{v}(wu)$
and
$LP(uvw)=w_{uv}\cdot e.lp_{w}(uv)+w_{vw}\cdot e.lp_{u}(vw)+w_{wu}\cdot e.lp_{v}(wu).$
Therefore, the function $\alpha LP(uvw)-ALG(uvw)$ is linear w.r.p. $(w_{uv},w_{vw},w_{wu})$.
Since the set of $(w_{uv},w_{vw},w_{wu})$ that satisfies the triangle inequality forms a convex cone generated by $(1,1,0),\,(1,0,1),\,(0,1,1)$, the function value is nonnegative for all such $(w_{uv},w_{vw},w_{wu})$ if and only if the value is nonnegative for $(w_{uv},w_{vw},w_{wu})\in\{(1,1,0),\,(1,0,1),\,(0,1,1)\}$. ∎
This lemma implies that the algorithm achieves an approximation factor of $\alpha$ if all of the following inequalities are satisfied for every possible configuration on the triangle $uvw$:
$e.cost_{w}(uv)+e.cost_{u}(vw)$ $\leq\alpha\cdot(e.lp_{w}(uv)+e.lp_{u}(vw)),$
$e.cost_{w}(uv)+e.cost_{v}(wu)$ $\leq\alpha\cdot(e.lp_{w}(uv)+e.lp_{v}(wu)),$
$e.cost_{u}(vw)+e.cost_{v}(wu)$ $\leq\alpha\cdot(e.lp_{u}(vw)+e.lp_{v}(wu)).$
We obtain a lower bound on the approximation factor of LP-Pivot by verifying the feasibility of rounding functions that satisfy the above inequalities. To this end, we analyze several configurations of LP values and edge signs on triangle $uvw$.
In Theorems 1 and 3, the notation ‘$(a,b,c)$ with $(s_{1},s_{2},s_{3})$’ denotes $(x_{uv},x_{vw},x_{wu})=(a,b,c)$, where each edge sign is given by $s_{1}$, $s_{2}$, and $s_{3}$, respectively.
Theorem 1.
The lower bound on the approximation factor of LP-Pivot in pseudometric-weighted correlation clustering is $10/3$.
Proof.
Suppose that LP-Pivot achieves an expected $\alpha$-approximation.
We begin by considering the configuration $(x,1-x,1)$ with signs (+,+,)(+,+,-)( + , + , - ) or (+,,)(+,-,-)( + , - , - ) and weights $(1,0,1)$. Then:
$\alpha(1-f^{+}(1-x))x$ $\geq(2-f^{+}(x))(1-f^{+}(1-x)),$ (14)
$\alpha(1-f^{-}(1-x))x$ $\geq(2-f^{+}(x))(1-f^{-}(1-x)),$ (15)
which implies:
$f^{+}(x)\leq 2-\alpha x\quad\implies\quad f^{+}(1-x)=f^{-}(1-x)=1.$ (16)
Next, consider the configuration $(0,x,x)$ with signs (+,+,+)(+,+,+)( + , + , + ) and weights $(1,1,0)$:
$\alpha x\geq 3f^{+}(x)-2f^{+}(x)^{2},$ (17)
which yields the bound $f^{+}(x)\leq\alpha x$. Therefore, for $x\in[0,1/\alpha]$, we have
$f^{+}(x)\leq\alpha x\leq 2-\alpha x,\quad f^{+}(1-x)=1.$
Now consider the configuration $(x,x,2x)$ with signs (+,+,+)(+,+,+)( + , + , + ) and weights $(1,1,0)$:
$\alpha x(1-f^{+}(x)f^{+}(2x))\geq f^{+}(x)+f^{+}(2x)-2f^{+}(x)f^{+}(2x).$ (18)
Rewriting the inequality, we obtain:
$(1-(2-\alpha x)f^{+}(x))f^{+}(2x)\leq\alpha x-f^{+}(x).$ (19)
Observe that for $x\in[0,1/\alpha)$ where $f^{+}(x)<1$, the term
$1-(2-\alpha x)f^{+}(x)\geq 1-2f^{+}(x)+f^{+}(x)^{2}\geq 0$
is strictly positive. Hence:
$f^{+}(2x)\leq\frac{\alpha x-f^{+}(x)}{1-(2-\alpha x)f^{+}(x)}=\alpha x-\frac{(1-\alpha x)^{2}f^{+}(x)}{1-(2-\alpha x)f^{+}(x)}\leq\alpha x,$ (20)
for $x\in[0,1/\alpha)$.
Combining the above, we conclude:
$f^{+}(x)<\frac{\alpha}{2}x\quad\text{for }x\in[0,2/\alpha).$
Set $x=\frac{4}{3\alpha}$. Then from (16) and (20), we have:
$f^{+}\left(\frac{4}{3\alpha}\right)\leq\frac{2}{3}=2-\alpha x,\quad\text{so}\quad f^{+}\left(1-\frac{4}{3\alpha}\right)=1.$
Since $f^{+}(x)<1$ for $x\in[0,2/\alpha)$, this implies:
$\frac{2}{\alpha}\leq 1-\frac{4}{3\alpha}\quad\Longrightarrow\quad\alpha\geq\frac{10}{3}.$
Hence, the approximation factor must be at least $10/3$. ∎
Conversely, there exist rounding functions $f^{+},\,f^{-}$ making the approximation factor of LP-Pivot by $10/3$, providing that the lower bound above is tight.
Theorem 2.
The LP-Pivot algorithm with the rounding function defined in equation 11 yields a $10/3$-approximation algorithm for pseudometric-weighted CC. The proof is provided in the Appendix.
5.2 Chromatic Correlation Clustering
We analyze the performance of the LP-CCC algorithm. This algorithm begins by assigning each vertex to its majority color based on the LP solution, followed by a pivot-based clustering routine.
Due to the strict majority condition, any edge not included in $\biguplus E_{c}$ must have an LP value of at least $1/2$. Thus, the cost incurred by such edges is at most twice their LP contribution XiuHTCH22 .
Within each color class $S_{c}$, corresponding to color $c$, we follow an analysis similar to that of Chawla et al. chawla2015near : edges of color $c$ are treated as positive edges ($E^{+}$), edges of the adversarial color $\gamma$ as negative edges ($E^{-}$), and all other edges as neutral ($E^{\circ}$).
For positive and negative edges, the definitions of $e.cost$ and $e.lp$ remain consistent with those in chawla2015near . The other three cases, particularly those involving neutral edges, require more careful treatment.
Consider a negative edge $uv\in E^{-}$: the LP value is
$e.lp_{w}(u,v)=\sum_{c^{\prime}\in L}(1-x_{uv}^{c^{\prime}})\geq 1-x_{uv}^{c}.$
For a neutral edge $uv\in E^{\circ}$, the expected cost arises from the event that $u$ and $v$ are not separated by $w$, i.e., at least one of them shares a cluster with $w$. The expected cost is thus given by the probability that $u$ and $v$ are not simultaneously separated from $w$.
The LP contribution in this case is the product of this probability with $x_{uv}^{\phi(uv)}$, where $\phi(uv)\neq c$ is the color of edge $uv$ in the input. While $x_{uv}^{\phi(uv)}$ is not tied to color $c$, we can still bound it below using $x_{uv}^{c},\,x_{vw}^{c},\,x_{wu}^{c}$ due to LP constraints XiuHTCH22 :
$x_{uv}^{\phi(uv)}$ $\geq\max\{x_{u}^{\phi(uv)},x_{v}^{\phi(uv)}\}$ (5)
$\geq\max\left\{\frac{1}{2},1-x_{u}^{c},1-x_{v}^{c}\right\}$ (7, 8)
$\geq\max\left\{\frac{1}{2},1-x_{uv}^{c},1-x_{vw}^{c},1-x_{wu}^{c}\right\}.$ (5)
Summarizing the results, we express the expected cost and lower bound on the LP value for a fixed pivot $w$ as follows:
$e.cost_{w}(u,v)=\begin{cases}p_{uw}(1-p_{vw})+(1-p_{uw})p_{vw},&uv\in E^{+};\\ (1-p_{uw})(1-p_{vw}),&uv\in E^{-};\\ 1-p_{uw}p_{vw},&uv\in E^{\circ};\end{cases}$ (21)
$e.lp_{w}(u,v)\geq\begin{cases}(1-p_{uw}p_{vw})x_{uv}^{c},&uv\in E^{+};\\ (1-p_{uw}p_{vw})(1-x_{uv}^{c}),&uv\in E^{-};\\ (1-p_{uw}p_{vw})\max\left\{\frac{1}{2},1-x_{uv}^{c},1-x_{vw}^{c},1-x_{wu}^{c}\right\},&uv\in E^{\circ}.\end{cases}$ (22)
These formulations are central to the analysis. Since $\alpha\cdot LP-ALG$ is always at least the expression obtained from the LP lower bound, we can prove that this bound is nonnegative.
As in Section 5.1, the algorithm achieves an $\alpha$-approximation if the following inequality holds for all triangles $uvw$:
$e.cost_{w}(uv)+e.cost_{u}(vw)+e.cost_{v}(wu)\leq\alpha\cdot\left(e.lp_{w}(uv)+e.lp_{u}(vw)+e.lp_{v}(wu)\right).$
This inequality leads to the following result on the approximation guarantee for LP-CCC:
Theorem 3.
The approximation factor of LP-CCC for CCC is bigger than $2.11$.
Proof.
Suppose LP-CCC is an expected $\alpha$-approximation algorithm, where $42\leq\alpha<$ 4.
Plugging $(\frac{1}{2},\frac{1}{2},1)$ with (,,)($\circ$,$\circ$,-)( ∘ , ∘ , - ), we obtain:
$\alpha\left(1-f^{\circ}\left(\frac{1}{2}\right)\right)\geq 2\left(1-f^{\circ}\left(\frac{1}{2}\right)\right)+\left(1-f^{\circ}\left(\frac{1}{2}\right)\right)^{2},$ (23)
which implies:
$f^{\circ}\left(\frac{1}{2}\right)\geq 3-\alpha.$ (24)
Next, plugging $\left(\frac{1}{4},\frac{1}{4},\frac{1}{2}\right)$ with (+,+,)(+,+$\circ$)( + , + , ∘ ) yields:
$\frac{1-f^{+}\left(\frac{1}{4}\right)^{2}+2\left(f^{+}\left(\frac{1}{4}\right)+f^{\circ}\left(\frac{1}{2}\right)-2f^{+}\left(\frac{1}{4}\right)f^{\circ}\left(\frac{1}{2}\right)\right)}{\frac{3}{4}\left(1-f^{+}\left(\frac{1}{4}\right)^{2}\right)+\frac{1}{2}\left(1-f^{+}\left(\frac{1}{4}\right)f^{\circ}\left(\frac{1}{2}\right)\right)}\leq\alpha,$ (25)
which can be rewritten as:
$\left(2+\left(\frac{\alpha}{2}-4\right)f^{+}\left(\frac{1}{4}\right)\right)f^{\circ}\left(\frac{1}{2}\right)\leq\frac{\alpha}{4}\left(5-3f^{+}\left(\frac{1}{4}\right)^{2}\right)+\left(f^{+}\left(\frac{1}{4}\right)-1\right)^{2}-2.$ (26)
Using the known bound from chawla2015near , plugging $(0,x,x)$ with (+,+,+)(+,+,+)( + , + , + ) yields:
$f^{+}(x)\leq 1-\sqrt{1-\alpha x},$ (27)
which implies:
$\left(2+\left(\frac{\alpha}{2}-4\right)f^{+}\left(\frac{1}{4}\right)\right)$ $\geq\left(2+\left(\frac{\alpha}{2}-4\right)\left(1-\sqrt{1-\frac{\alpha}{4}}\right)\right)$
$=2\sqrt{1-\frac{\alpha}{4}}\left(\left(\sqrt{1-\frac{\alpha}{4}}-\frac{1}{2}\right)^{2}+\frac{3}{4}\right)>0,$
for any $\alpha<4$.
Therefore, we obtain the following upper bound on $f^{\circ}\left(\frac{1}{2}\right)$:
$f^{\circ}\left(\frac{1}{2}\right)\leq\frac{\frac{\alpha}{4}\left(5-3f^{+}\left(\frac{1}{4}\right)^{2}\right)+\left(f^{+}\left(\frac{1}{4}\right)-1\right)^{2}-2}{2+\left(\frac{\alpha}{2}-4\right)f^{+}\left(\frac{1}{4}\right)}.$ (28)
Define
$g(x,y):=\frac{\frac{y}{4}\left(5-3x^{2}\right)+\left(x-1\right)^{2}-2}{2+\left(\frac{y}{2}-4\right)x}$
for $x\in\left[0,1-\sqrt{1-\frac{y}{4}}\right]$ and $y\in[2,4)$, to examine the maximum possible value of $f^{\circ}\left(\frac{1}{2}\right)$ for a given $\alpha$.
The partial derivative of $g(x,y)$ with respect to $x$ is:
$\frac{\partial}{\partial x}g(x,y)=\frac{3y-4}{2(8-y)}+\frac{5y^{3}-84y^{2}+368y-448}{2(8-y)(4-8x+xy)^{2}}.$ (29)
Within the given region, $\frac{\partial}{\partial x}g(x,y)$ is nonzero, as shown in Figure 2. Moreover, evaluating at $x=0$ gives:
$\left.\frac{\partial}{\partial x}g(x,y)\right|_{x=0}=\frac{5y^{3}-84y^{2}+416y-512}{32(8-y)}=-\frac{5y^{2}-44y+64}{32}\geq\frac{1}{8}.$ (30)
Figure 2: Comparison between the domain of $g$ and the curve where $\frac{\partial}{\partial x}g(x,y)=0.$
Therefore, for any $y_{0}\in[2,4)$, the function $g(x,y_{0})$ is strictly increasing within the region. Consequently,
$g(x,y_{0})\leq g\left(1-\sqrt{1-\frac{y_{0}}{4}},y_{0}\right),$
which corresponds to the parabolic boundary of the region depicted in Figure 2.
Parameterizing this curve as $(1-t,4(1-t^{2}))$ for $t\in\left(0,\frac{1}{\sqrt{2}}\right]$, the value of $g$ along the curve is:
$g(1-t,4(1-t^{2}))$ $=\frac{(1-t^{2})(2+6t-3t^{2})+t^{2}-2}{2-2(t^{2}+1)(1-t)}$ (31)
$=\frac{3t^{3}-6t^{2}-4t+6}{2(t^{2}-t+1)}.$ (32)
This is the upper bound for $f^{\circ}\left(\frac{1}{2}\right)$ when $\alpha=4(1-t^{2})$.
On the other hand, from (24), the lower bound is $13-\alpha=3-4(1-t^{2})=4t^{2}-$ 1.
Therefore, the inequality
$62(t^{2}-t+1)(4t^{2}-1)\leq 3t^{3}-6t^{2}-4t+$ 6 (33)
must hold.
Define:
$h(t):=2(t^{2}-t+1)(4t^{2}-1)-(3t^{3}-6t^{2}-4t+6)=8t^{4}-11t^{3}+12t^{2}+6t-8.$
Since $h^{\prime\prime}(t)=96t^{2}-66t+24>0$ and $h^{\prime}(0)=6>0$, $h$ is convex and increasing for $t\geq 0$.
Finally, since $h\left(\sqrt{1-\frac{2.11}{4}}\right)>0$, we conclude that $h(t)>0$ when $t\geq\sqrt{1-\frac{2.11}{4}}$, implying that the lower bound for $f^{\circ}\left(\frac{1}{2}\right)$ exceeds the upper bound when $\alpha\leq 2.11$.
Therefore, the approximation factor of LP-CCC for CCC must be greater than $2.11$. ∎
Analogous to the classical CC setting—where the lower bound and the approximation ratio of LP-Pivot differ by less than $0.04$ chawla2015near —augmenting the LP rounding with a suitable $f^{\circ}$ yields the following:
Theorem 4.
LP-CCC, using the rounding functions in (12) and (13), achieves a $2.15$-approximation for Chromatic Correlation Clustering.
5.3 Proof for Theorem 4
Unlike the pseudometric-weighted case, the variables $x_{vw}^{c}$ and $x_{wu}^{c}$ can multiplicatively affect the value of $e.lp_{w}(u,v)$ in (22). Therefore, Lemma 5 does not apply, and we must instead verify the inequality $\alpha\cdot LP-ALG\geq 0$ directly.
For sign configurations that do not involve a $\circ$’ (i.e., all edges have $+$ or -- signs), the inequality $\alpha\cdot LP-ALG\geq 0$ with $\alpha=2.06-\varepsilon$ is already established in chawla2015near . Thus, it suffices to check the remaining $6$ configurations involving at least one $\circ$’ sign.
As before, define $(x,y,z):=(x_{uv},x_{vw},x_{wu})$ and $(p_{x},p_{y},p_{z}):=(p_{uv},p_{vw},p_{wu})$.
5.3.1 (,,)($\circ$,$\circ$,$\circ$)( ∘ , ∘ , ∘ ) Triangles
$2\cdot e.lp_{w}(u,v)\geq 2(1-p_{uw}p_{vw})\cdot\frac{1}{2}=(1-p_{uw}p_{vw})=e.cost_{w}(u,v).$
Since the same inequality holds for all permutations of the endpoints, we conclude:
$2\cdot LP-ALG\geq 0$
5.3.2 (,,+)($\circ$,$\circ$,+)( ∘ , ∘ , + ) Triangles
If $z\geq\frac{1}{2}$,
$2\cdot e.lp_{v}(w,u)-e.cost_{v}(w,u)\geq(1-p_{x}p_{y})-(p_{x}+p_{y}-2p_{x}p_{y})=(1-p_{x})(1-p_{y})\geq 0,$
and by the same reasoning as in Section 5.3.1, we also have
$2\cdot e.lp_{w}(u,v)\geq e.cost_{w}(u,v),\quad 2\cdot e.lp_{u}(v,w)\geq e.cost_{u}(v,w).$
Hence, $2\cdot LP-ALG\geq 0$.
Otherwise, i.e., if $1-2z\geq 0$:
$2\cdot LP-ALG$ $=2\cdot\left((2-(p_{x}+p_{y})p_{z})\max\left\{\frac{1}{2},1-x,1-y,1-z\right\}+(1-p_{x}p_{y})z\right)$
$\quad-\left(1-p_{y}p_{z}+1-p_{x}p_{z}+p_{x}+p_{y}-2p_{x}p_{y}\right)$
$\geq 2\cdot\left((2-(p_{x}+p_{y})p_{z})(1-z)+(1-p_{x}p_{y})z\right)$
$\quad-\left(2-(p_{x}+p_{y})p_{z}+(1-p_{x}p_{y})-(1-p_{x})(1-p_{y})\right)$
$=(1-2z)(2-(p_{x}+p_{y})p_{z})+(1-2z)(p_{x}p_{y}-1)+(1-p_{x})(1-p_{y})$
$=(1-2z)\left(2-(p_{x}+p_{y})p_{z}+p_{x}p_{y}-1\right)+(1-p_{x})(1-p_{y})$
$\geq(1-2z)\left(2-p_{x}-p_{y}+p_{x}p_{y}-1\right)+(1-p_{x})(1-p_{y})$
$=(1-2z)(1-p_{x})(1-p_{y})+(1-p_{x})(1-p_{y})\geq 0.$
5.3.3 (,,)($\circ$,$\circ$,-)( ∘ , ∘ , - ) Triangles
If $z\leq 1-\frac{1}{\alpha}$, then as in Section 5.3.2, we have $\alpha\cdot LP-ALG\geq 0$.
Otherwise, since $21-z\leq 1/\alpha<1/$ 2,
$\alpha\cdot LP-ALG$
$=\alpha\cdot\left((2-(p_{x}+p_{y})z)\max\left\{\frac{1}{2},1-x,1-y,1-z\right\}+(1-p_{x}p_{y})(1-z)\right)$
$\quad-\left(1-p_{y}z+1-p_{x}z+(1-p_{x})(1-p_{y})\right)$
$=\alpha(2-(p_{x}+p_{y})z)\max\left\{\frac{1}{2},1-x,1-y\right\}+\alpha(1-p_{x}p_{y})(1-z)$
$\quad-\left(2-(p_{x}+p_{y})z+(1-p_{x})(1-p_{y})\right).$
Fixing the values of $x$ and $y$, the function is affine with respect to $z$, and its slope is given by
$-\alpha(p_{x}+p_{y})\max\left\{\frac{1}{2},1-x,1-y\right\}-\alpha(1-p_{x}p_{y})+(p_{x}+p_{y})$
$\leq\left(-\frac{\alpha}{2}+1\right)(p_{x}+p_{y})-\alpha(1-p_{x}p_{y})\leq 0.$
Therefore, the function is minimized when $z$ is maximized.
Now, w.l.o.g., assume $x\leq y$, and consider the following three cases:
  1. 1.
    $x\geq 1/2$
    $p_{x},p_{y}\geq 0.85$ as $x,y\geq 1/2$. Applying $z=1$,
    $\alpha\cdot LP-ALG$
    $\geq\alpha(2-p_{x}-p_{y})\max\left\{\frac{1}{2},1-x,1-y\right\}-(2-p_{x}-p_{y})-(1-p_{x})(1-p_{y})$
    $=\left(\frac{\alpha}{2}-1\right)(2-p_{x}-p_{y})-(1-p_{x})(1-p_{y})$
    $\geq\left(\frac{\alpha}{2}-1\right)(2-p_{x}-p_{y})-\frac{1}{4}(2-p_{x}-p_{y})^{2}$
    $=\frac{1}{4}(2-(p_{x}+p_{y}))(2\alpha-6+(p_{x}+p_{y})).$
    Since $6-2\alpha=1.7\leq p_{x}+p_{y}\leq 2$, $\alpha\cdot LP-ALG\geq 0$.
  2. 2.
    $x<1/2,\,x+y\geq 1$
    Applying $z=1$ again,
    $\alpha\cdot LP-ALG$
    $\geq\alpha(2-p_{x}-p_{y})\max\left\{\frac{1}{2},1-x,1-y\right\}-(2-p_{x}-p_{y})-(1-p_{x})(1-p_{y})$
    $=(\alpha(1-x)-1)(2-p_{x}-p_{y})-(1-p_{x})(1-p_{y}).$
    Note that the function is affine with respect to $p_{y}$; therefore, the function is minimized when $p_{y}$ reaches the boundary, i.e., $y=1$ or $y=1-x$.
    If $y=1$, then $p_{y}=1$, hence
    $(\alpha(1-x)-1)(2-p_{x}-p_{y})-(1-p_{x})(1-p_{y})$
    $=(\alpha(1-x)-1)(1-p_{x})$
    $\geq\left(\frac{\alpha}{2}-1\right)(1-p_{x})\geq 0.$
    On the other hand, if $y=1-x$, then $p_{y}=\frac{3}{10}y+\frac{7}{10}=1-\frac{3}{10}x$ as $y>1/2$, hence
    $(\alpha(1-x)-1)(2-p_{x}-p_{y})-(1-p_{x})(1-p_{y})$
    $=(\alpha(1-x)-1)\left(1-\frac{17}{10}x+\frac{3}{10}x\right)-\frac{3}{10}\left(1-\frac{17}{10}x\right)x$
    $=3.52\cdot\left(x-\frac{1}{2}\right)\left(x-\frac{115}{176}\right)\geq 0$
    as $x<1/2$ and $p_{x}=\frac{17}{10}x$.
  3. 3.
    $x+y<1$
    Applying $z=x+y$,
    $\alpha\cdot LP-ALG$
    $=\alpha(2-(p_{x}+p_{y})(x+y))\max\left\{\frac{1}{2},1-x,1-y\right\}+\alpha(1-p_{x}p_{y})(1-x-y)$
    $\quad-\left(2-(p_{x}+p_{y})(x+y)+(1-p_{x})(1-p_{y})\right)$
    $=(\alpha(1-x)-1)(2-(p_{x}+p_{y})(x+y))+\alpha(1-p_{x}p_{y})(1-x-y)-(1-p_{x})(1-p_{y})$
    $\geq 0.3\cdot(\alpha(1-x)-1)+\alpha(1-p_{x}p_{y})(1-x-y)-(1-p_{x})(1-p_{y}).$
    For the last inequality, since $p_{x}\leq 1.7x$ and $p_{y}\leq 1.7y$, $2-(p_{x}+p_{y})(x+y)\geq 2-1.7(x+y)^{2}\geq 0.3$.
    Differentiating the function by $y$,
    $-\alpha p_{x}\frac{dp_{y}}{dy}\cdot(1-x-y)-\alpha(1-p_{x}p_{y})+\frac{dp_{y}}{dy}\cdot(1-p_{x})$
    $\leq-\alpha p_{x}\frac{dp_{y}}{dy}\cdot(1-x-y)-\alpha(1-p_{x})+\frac{dp_{y}}{dy}\cdot(1-p_{x})<0$
    since $\frac{dp_{y}}{dy}\leq 1.7<\alpha$. Hence, the function is minimized when $y$ is maximized.
    Applying $y=1-x$, since $y=1-x>1/2$, $p_{y}=\frac{3}{10}y+\frac{7}{10}=1-\frac{3}{10}x$ and thus
    $0.3\cdot(\alpha(1-x)-1)+\alpha(1-p_{x}p_{y})(1-x-y)-(1-p_{x})(1-p_{y})$
    $=0.3\cdot(\alpha(1-x)-1)-\left(1-\frac{17}{10}x\right)\cdot\frac{3}{10}x$
    $=0.51\left(x-\frac{1}{2}\right)\left(x-\frac{23}{17}\right)\geq 0$
    since $x<1/2$.
5.3.4 (,+,+)($\circ$,+,+)( ∘ , + , + ) Triangles
For Sections 5.3.4 and 5.3.5, we utilize the following lemma, which serves as a weaker version of Lemma 5:
Lemma 3.
Let the sign configuration be such that the edge $uv$ is the only edge labeled with a $\circ$’ among the three edges of the triangle. Suppose the rounding function $f^{\circ}$ is piecewise affine. Then $\mathcal{C}\geq 0$ for all possible configurations of $(x_{uv},x_{vw},x_{wu})$ if the inequality holds in each of the following cases:
  1. 1.
    The triangle inequality is tight for $(x_{uv},x_{vw},x_{wu})$;
  2. 2.
    $x_{uv}$ lies at an endpoint of a domain interval for some piece of the function $f^{\circ}$;
  3. 3.
    $x_{uv}=\min\left\{\frac{1}{2},x_{vw},x_{wu}\right\}$.
Proof.
Fix the values of $x_{vw}$ and $x_{wu}$, and treat $\mathcal{C}$ as a function of $x_{uv}$ and $p_{uv}$. Within this function, the term involving $x_{uv}$ appears only in
$e.lp_{w}(uv)=(1-p_{vw}p_{wu})\cdot\max\left\{\frac{1}{2},1-x_{uv},1-x_{vw},1-x_{wu}\right\},$
which is piecewise affine with two domain intervals:
$\left[0,\min\left\{\frac{1}{2},x_{vw},x_{wu}\right\}\right]\quad\text{and}\quad\left[\min\left\{\frac{1}{2},x_{vw},x_{wu}\right\},1\right].$
The remaining terms in $\mathcal{C}$ are affine in $p_{uv}$.
Hence, $\mathcal{C}$ is a piecewise affine function defined over two domains in the $(x_{uv},p_{uv})$ space:
$\left[0,\min\left\{\frac{1}{2},x_{vw},x_{wu}\right\}\right]\times[0,1]\quad\text{and}\quad\left[\min\left\{\frac{1}{2},x_{vw},x_{wu}\right\},1\right]\times[0,1].$
Since $f^{\circ}$ is piecewise affine, substituting $p_{uv}=f^{\circ}(x_{uv})$ yields a piecewise affine function in $x_{uv}$. The breakpoints of this composition occur at:
  1. 1.
    the endpoints of the domain intervals of $f^{\circ}$, and
  2. 2.
    the point $x_{uv}=\min\left\{\frac{1}{2},x_{vw},x_{wu}\right\}$ where the domain of the piecewise max function changes.
Therefore, the minimum of $\mathcal{C}$ occurs either at domain boundaries or at breakpoints—precisely the three cases specified in the lemma. ∎
Moreover, in this particular case, an even stronger argument can be made: Examining the full formula of $\mathcal{C}=\alpha\cdot LP-ALG$ with this sign configuration,
$\alpha\cdot LP-ALG$
$=\alpha\cdot\left((1-p_{y}p_{z})\max\left\{\frac{1}{2},1-x,1-y,1-z\right\}+(1-p_{x}p_{z})y+(1-p_{x}p_{y})z\right)$
$\quad-\left(1-p_{y}p_{z}+p_{x}+p_{z}-2p_{x}p_{z}+p_{x}+p_{y}-2p_{x}p_{y}\right)$
$=\alpha(1-p_{y}p_{z})\max\left\{\frac{1}{2},1-x,1-y,1-z\right\}+\left(-\alpha(yp_{z}+p_{y}z)-2+2(p_{y}+p_{z})\right)p_{x}$
$\quad+\alpha(y+z)-(1+p_{y}+p_{z}-p_{y}p_{z})$
The first term is a monotonically decreasing function of $x$; the second term is a monotonically decreasing function of $p_{x}$ since $-\alpha(yp_{z}+p_{y}z)-2+2(p_{y}+p_{z})\leq 0$, as shown in the following Lemma:
Lemma 4.
For $f^{+}$ given as equation 12 and $\alpha\in\left[\frac{1}{0.5095},\frac{2}{3\cdot 0.5095-0.19}\right]\approx[1.963,2.988]$,
$\alpha(xf^{+}(y)+f^{+}(x)y)+2-2(f^{+}(x)+f^{+}(y))\geq 0$
in $[0,1]^{2}$.
Proof.
w.l.o.g. $x\leq y$. Consider the $2$ possible cases, based on the value of $y$.
  1. 1.
    $y\geq 0.5095$
    In this case, $f^{+}(y)=1$ and thus
    $\alpha(xf^{+}(y)+f^{+}(x)y)+2-2(f^{+}(x)+f^{+}(y))$
    $=\alpha(x+f^{+}(x)y)-2f^{+}(x),$
    which is an increasing function of $y$.
    If $x\geq 0.5095$, setting $y=x$,
    $\alpha(x+f^{+}(x)y)-2f^{+}(x)\geq 2\alpha x-2\geq 0$
    since $x>1/2$. Otherwise, setting $y=0.5095$,
    $\alpha(x+f^{+}(x)y)-2f^{+}(x)\geq(\alpha\cdot 0.5095-2)f^{+}(x)+\alpha x$
    is concave of $x$ since $\alpha\cdot 0.5095-2<0$ and $f^{+}$ is convex in the range. Hence, this is minimized at either $x=0$ or $x=0.5095$, where the former case results in 00 and the latter case results in $2\alpha\cdot 0.5095-2\geq 0$.
  2. 2.
    $y<0.5095$
    As a function of $y$,
    $\alpha(xf^{+}(y)+f^{+}(x)y)+2-2(f^{+}(x)+f^{+}(y))$
    $=(\alpha x-2)f^{+}(y)+\alpha f^{+}(x)y+2-2f^{+}(x)$
    is concave since $\alpha x-2\leq\alpha\cdot 0.5095-2<0$ and $f^{+}$ is convex in the range. Hence, this is minimized at either $y=x$ or $y=0.5095$, where the latter case reduce to Case 1, and the former case results in
    $(\alpha x-2)f^{+}(y)+\alpha f^{+}(x)y+2-2f^{+}(x)$
    $=2\left((\alpha x-2)f^{+}(x)+1\right),$
    which is $2\geq 0$ if $x<0.19$, or
    $2\left((\alpha x-2)\left(\frac{x-0.19}{0.5095-0.19}\right)^{2}+1\right)$
    otherwise. This is a decreasing function when $x\in\left[0,\frac{0.19+4/\alpha}{3}\right]$, which contains $[0,0.5095]$. Therefore, this is minimized when $x=0.5095$, which also reduce to Case 1.
This completes the proof of the lemma. ∎
Therefore, as $p_{x}$ is a monotonically increasing function of $x$, the function is minimized when $x$ is maximized.
Now, w.l.o.g. $y\leq z$ and consider all $2$ possible cases:
  1. 1.
    $y+z\geq 1$
    Applying $x=1$,
    $\alpha\cdot LP-ALG$
    $\geq\alpha\cdot\left((1-p_{y}p_{z})\max\left\{\frac{1}{2},1-y,1-z\right\}+(1-p_{y})z+(1-p_{z})y\right)$
    $\quad-(3-p_{y}p_{z}-p_{z}-p_{y})$
    $\geq\alpha\cdot\left(\frac{1}{2}(1-p_{y}p_{z})+(1-p_{z})y+(1-p_{y})z\right)-(3-p_{y}p_{z}-p_{z}-p_{y})$
    $\geq\alpha\cdot\left(\frac{1}{2}(1-p_{y}p_{z})+(1-p_{y})\frac{p_{z}}{2}+(1-p_{z})\frac{p_{y}}{2}\right)-(3-p_{y}p_{z}-p_{z}-p_{y})$
    $=2.225\left(\left(\frac{6}{89}\right)^{2}-\left(p_{y}-\frac{83}{89}\right)\left(p_{z}-\frac{83}{89}\right)\right).$
    Since $z\geq 1/2$, $p_{z}>0.94>\frac{83}{89}$, thus the function is decreasing with respect to $p_{y}$. The minimum value of the function is therefore 00 when $p_{y}=p_{z}=1$.
  2. 2.
    $y+z<1$
    We use the similar argument with of those conducted by Chawla et al.chawla2015near on the (+,+,)(+,+,-)( + , + , - ) case. For simplicity, let $a=0.19,\,b=0.5095$.
    If $y<a$, $p_{y}=0$, hence
    $\alpha\cdot LP-ALG$
    $=\alpha\cdot\left((1-y)+(1-p_{x}p_{z})y+z\right)-\left(1+2p_{x}+p_{z}-2p_{x}p_{z}\right)$
    $=-\alpha p_{x}p_{z}y+\alpha(1+z)-\left(1+2p_{x}+p_{z}-2p_{x}p_{z}\right).$
    This is a non-increasing function of $y$, and thus it is sufficient to check on $y=a$.
    If $z>b$, $p_{z}=1$, hence
    $\alpha\cdot LP-ALG$
    $=\alpha\cdot\left((1-p_{y})(1-y)+(1-p_{x})y+(1-p_{x}p_{y})z\right)-\left(2-2p_{x}p_{y}\right)$
    $=\alpha(1-p_{x}p_{y})z+\alpha((1-p_{y})(1-y)+(1-p_{x})y)-2(1-p_{x}p_{y}).$
    This is a non-decreasing function of $z$, and thus it is sufficient to check on $z=b$.
    Now the problem reduces to only $a\leq y\leq z\leq b$ case, where the formula for $p_{y}$ and $p_{z}$ are both quadratic.
    We now show that the function is minimized when $y=z$ as well as $x=y+z$. Consider the parametrization $(x,y,z)=(x,t-c,t+c)$, where $t\in[a,1/2],\,c\in[0,\min\{t-a,b-t\}],$ and $x=2t$ since $\alpha\cdot LP-ALG$ is a monotonically increasing function of $x$. Then
    $\alpha\cdot LP-ALG$
    $=\alpha\cdot\left((1-p_{y}p_{z})(1-t)+(1-p_{x}p_{z})(t-c)+(1-p_{x}p_{y})(t+c)\right)$
    $\quad-\left(1+2p_{x}+p_{y}+p_{z}-p_{y}p_{z}-2p_{x}p_{z}-2p_{x}p_{y}\right)$
    $\quad+\alpha(1-p_{y})c.$
    Fixing the value of $t$, the sum of first $2$ terms is a negative even quartic function of $c$, where the negativity is obtained from $-\alpha(1-t)+1\leq-\alpha(1-b)+1<0$. Thus, we can write the function as
    $P(t)c^{4}+Q(t)c^{2}+\alpha(1-p_{y})c+R(t),$
    where $P(t)=\frac{-\alpha(1-t)+1}{(b-a)^{4}}<0$. Since $\alpha(1-p_{y})c\geq 0$ if $c\geq 0$, it is sufficient to show that $Q\geq 0$ and $\sqrt{\frac{Q}{-P}}\geq\min\{t-a,b-t\}$ to prove that the function is minimized when $c=0$.
    Calculating the $Q$, this results in
    $Q(t)$ $=\frac{2}{(b-a)^{2}}\cdot\left((\alpha(1-t)-1)\left(\frac{t-a}{b-a}\right)^{2}+\alpha p_{x}(t-2a)-(1-2p_{x})\right)$
    $=-P(t)\cdot(t-a)^{2}+\frac{2}{(b-a)^{2}}\cdot\left(-1+(\alpha t+2-2\alpha a)p_{x}\right)$
    $\geq-P(t)\cdot(t-a)^{2},$
    where the final inequality is derived from
    $\frac{1}{\alpha t+2-2\alpha a}\leq\frac{1}{2-\alpha a}\approx 0.628<0.646=f^{\circ}(2\cdot 0.19)\leq p_{x}.$
    Hence, $Q\geq-P(t)\cdot(t-a)^{2}>0$ and therefore $\sqrt{\frac{Q}{-P}}\geq t-a\geq\min\{t-a,b-t\}$, which implies that $\alpha\cdot LP-ALG$ is minimized at $c=0$, i.e. at $(2t,t,t)$.
    Finally, plugging $(x,y,z)=(2t,t,t)$,
    $\alpha\cdot LP-ALG$
    $=\alpha\cdot\left((1-p_{y}^{2})(1-y)+2(1-p_{x}p_{y})y\right)-\left(1+2p_{x}+2p_{y}-p_{y}^{2}-4p_{x}p_{y}\right)$
    $=\left(\alpha(1-p_{y}^{2})(1-t)+2\alpha t-1-2p_{y}+p_{y}^{2}\right)-(2\alpha p_{y}t+2-4p_{y})p_{x}.$
    The top left region of Figure 1 shows the region $(2t,f^{\circ}(2t))$ where the formula above is negative when $f^{+}$ is given as equation 12. Since our proposed $f^{\circ}$ doesn’t intersect the region, $\alpha\cdot LP-ALG\geq 0$ if $y=z$, and thus $\alpha\cdot LP-ALG\geq 0$ whenever $a\leq y\leq z\leq b$.
5.3.5 (,+,)($\circ$,+,-)( ∘ , + , - ) Triangles
We would prove that the lower bound $\mathcal{C}^{\prime}$ of $\mathcal{C}$, which is defined as
$\alpha\cdot LP-ALG$
$=\alpha\cdot\left((1-p_{y}z)\max\left\{\frac{1}{2},1-x,1-y,1-z\right\}+(1-p_{x}p_{y})(1-z)+(1-p_{x}z)y\right)$
$\quad-\left(1-p_{y}z+p_{x}+z-2p_{x}z+(1-p_{x})(1-p_{y})\right)$
$\geq\alpha\cdot\left((1-p_{y}z)\max\left\{\frac{1}{2},1-y,1-z\right\}+(1-p_{x}p_{y})(1-z)+(1-p_{x}z)y\right)$
$\quad-\left(1-p_{y}z+p_{x}+z-2p_{x}z+(1-p_{x})(1-p_{y})\right)\triangleq\mathcal{C}^{\prime}$
is nonnegative.
Since this is now affine with respect to $p_{x}$, this further enhances the argument Lemma 3 to check only on the endpoint.
  1. 1.
    $y\geq 1/2,\,z\leq 1/2$
    As in 5.3.2 and 5.3.3, $2\cdot LP-ALG\geq 0$.
  2. 2.
    $y,z\geq 1/2$
    Similar with 5.3.4, we can deduce a stronger argument than of Lemma 3: Examining the full formula of $\mathcal{C}^{\prime}$,
    $\mathcal{C}^{\prime}$ $=\alpha\cdot\left((1-p_{y}z)\max\left\{\frac{1}{2},1-y,1-z\right\}+(1-p_{x}p_{y})(1-z)+(1-p_{x}z)y\right)$
    $\quad-\left(1-p_{y}z+p_{x}+z-2p_{x}z+(1-p_{x})(1-p_{y})\right)$
    $=(-\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z)p_{x}$
    $\quad+\alpha(1-z)+\alpha y+p_{y}z-z+p_{y}-2+\alpha(1-p_{y}z)\max\left\{\frac{1}{2},1-y,1-z\right\}.$
    The first term is a monotonically decreasing function of $p_{x}$ since
    $-\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z=(\alpha(p_{y}-y)+2)z-(\alpha+1)p_{y}$
    is an increasing function of $z$ since $p_{y}\geq y$ in the range,
    $-\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z\leq-\alpha y-p_{y}+2\leq-\frac{\alpha}{2}-f^{+}\left(\frac{1}{2}\right)+2\approx-0.016<0.$
    Hence, the function is minimized when $x$ is maximized. Setting $x=1$,
    $\mathcal{C}^{\prime}$ $\geq\alpha\cdot\left((1-p_{y}z)\cdot\frac{1}{2}+(1-p_{y})(1-z)+(1-z)y\right)-(2-p_{y}z-z)$
    $=\left(-\alpha+1-\alpha y+\left(\frac{\alpha}{2}+1\right)p_{y}\right)z+\alpha y-\alpha p_{y}+\frac{3}{2}\alpha-2.$
    Since $y\geq 1/2$ and $p_{y}\leq 1$,
    $-\alpha+1-\alpha y+\left(\frac{\alpha}{2}+1\right)p_{y}\leq-\alpha+2<0.$
    Hence, the function is minimized when $z$ is maximized. Setting $z=1$,
    $\mathcal{C}^{\prime}$ $\geq\frac{\alpha}{2}-1+\left(-\frac{\alpha}{2}+1\right)p_{y}$
    $=\left(\frac{\alpha}{2}-1\right)(1-p_{y})\geq 0.$
  3. 3.
    $y\leq z,\,y<1/2$
    $\mathcal{C}^{\prime}$ $=\alpha\cdot\left((1-p_{y}z)(1-y)+(1-p_{x}p_{y})(1-z)+(1-p_{x}z)y\right)$
    $\quad-\left(1-p_{y}z+p_{x}+z-2p_{x}z+(1-p_{x})(1-p_{y})\right)$
    $=(-\alpha(1+p_{y}-p_{y}y-p_{x}p_{y}+p_{x}y)+2p_{x}+p_{y}-1)z$
    $\quad+\alpha(2-p_{x}p_{y})-2+p_{y}-p_{x}p_{y}.$
    This is a decreasing function of $z$ since
    $-\alpha(1+p_{y}-p_{y}y-p_{x}p_{y}+p_{x}y)+2p_{x}+p_{y}-1$
    $=(2+\alpha p_{y}-\alpha y)p_{x}-\alpha(1+p_{y}-p_{y}y)+p_{y}-1$
    is an increasing function of $p_{x}$ since $2+\alpha p_{y}-\alpha y\geq 2-\frac{\alpha}{2}\geq 0$,
    $-\alpha(1+p_{y}-p_{y}y-p_{x}p_{y}+p_{x}y)+2p_{x}+p_{y}-1$
    $\leq-\alpha(1-p_{y}y+y)+1+p_{y}$
    $=-(1-p_{y})(1+\alpha y)-\alpha+2\leq-\alpha+2<0.$
    Hence, the function is minimized when $z$ is maximized. Setting $z=1$,
    $\mathcal{C}^{\prime}$ $\geq\alpha\cdot\left((1-p_{y})(1-y)+(1-p_{x})y\right)-\left(3-2p_{x}-2p_{y}+p_{x}p_{y}\right).$
    As mentioned above, it is sufficient to check on $x=1$ and $x=1-y$.
    1. (a)
      $x=1$
      $\mathcal{C}^{\prime}$ $=\alpha\cdot(1-p_{y})(1-y)-(1-p_{y})$
      $=(1-p_{y})(\alpha-1-\alpha y)$
      $\geq(1-p_{y})\left(\frac{\alpha}{2}-1\right)\geq 0.$
    2. (b)
      $x=1-y$
      Since $x=1-y>1/2$, $p_{x}=\frac{3}{10}x+\frac{7}{10}=1-\frac{3}{10}y$. Therefore,
      $\mathcal{C}^{\prime}$ $=\alpha\cdot\left((1-p_{y})(1-y)+\frac{3}{10}y^{2}\right)-\left(1+\frac{3}{10}y\right)(2-p_{y})+1.$
      If $y<0.19$,
      $\mathcal{C}^{\prime}$ $=\alpha\cdot\left(1-y+\frac{3}{10}y^{2}\right)-2\left(1+\frac{3}{10}y\right)+1$
      $=\frac{1}{200}(129y^{2}-550y+230)>0.65\geq 0.$
      Otherwise, i.e. $0.50950.19\leq y<1/2<$ 0.5095,
      $\mathcal{C}^{\prime}\approx 0.74331+2.39737y-19.7409y^{2}+24.0007y^{3},$
      whose minimum value in the region is approximately $7.236\times 10^{-6}\geq 0$ in $y\approx 0.4788$.
  4. 4.
    $z<y<1/2$
    $\mathcal{C}^{\prime}$ $=\alpha\cdot\left((1-p_{y}z)(1-z)+(1-p_{x}p_{y})(1-z)+(1-p_{x}z)y\right)$
    $\quad-\left(1-p_{y}z+p_{x}+z-2p_{x}z+(1-p_{x})(1-p_{y})\right)$
    $=\alpha p_{y}z^{2}+(-\alpha(2+p_{y}-p_{x}p_{y}+p_{x}y)+2p_{x}+p_{y}-1)z$
    $\quad+\alpha(2+y-p_{x}p_{y})-2+p_{y}-p_{x}p_{y}.$
    Differentiating by $z$,
    $1\displaystyle 2\alpha p_{y}z-\alpha(2+p_{y}-p_{x}p_{y}+p_{x}y)+2p_{x}+p_{y}-$ 1
    $\leq\alpha p_{y}-\alpha(2+p_{y}-p_{x}p_{y}+p_{x}y)+2p_{x}+p_{y}-1$
    $=(-2\alpha+2p_{x}+p_{y})-\alpha((1-p_{x})p_{y}+p_{x}y)-1\leq 0$
    since $z<1/2$ and $-2\alpha+2p_{x}+p_{y}\leq-2\alpha+3=-1.3<0$. Therefore, the function is minimized when $z$ is maximized.
    Plugging $z=y$, this now reduce to Case 3 with $z=y$ due to the continuity of $\mathcal{C}^{\prime}$.
5.3.6 (,,)($\circ$,-,-)( ∘ , - , - ) Triangles
As in 5.3.1, $2e.lp_{w}(u,v)\geq e.cost_{w}(u,v)$. For other two terms,
$e.lp_{u}(v,w)+e.lp_{v}(w,u)$ $\geq(1-p_{uv}z)(1-y)+(1-p_{uv}y)(1-z)$
$\geq(1-p_{uv})(1-y)+(1-p_{uv})(1-z)$
$=(1-p_{uv})(1-p_{vw})+(1-p_{uv})(1-p_{wu})$
$=e.cost_{v}(w,u)+e.cost_{u}(v,w).$
Hence, $2\cdot LP-ALG\geq 0$.
This completes our case analysis.
6 Conclusion
In this work, we studied two important variants of correlation clustering: pseudometric-weighted correlation clustering and chromatic correlation clustering (CCC). For both problems, we developed and analyzed specialized rounding functions that are essential for achieving improved approximation guarantees via the LP-Pivot algorithm.
For the pseudometric-weighted setting, we proposed a piecewise-linear rounding function tailored for the setting that achieves a $10/3$-approximation, and showed that no alternative function within our analytical framework can yield a better factor. For the chromatic correlation clustering (CCC) variant, we designed a distinct rounding function that respects the constraints imposed by color restrictions and achieves an approximation factor of $2.15$. The function is constructed using a piecewise-linear form and leverages a careful analysis of triple costs.
Overall, our work highlights the importance of designing principled and variant-specific rounding strategies to extend LP-based techniques to structured clustering problems, yielding strong theoretical guarantees.
Appendix A Proof for $6$-approximation on Pseudometric-weighted CC with LP-UMVD-Pivot
With the parameters $\alpha=\frac{1}{3}$ and $\beta=0$, the corresponding rounding functions are as follows:
$f^{+}(x)=f^{-}(x)=\begin{cases}0,&x<\frac{1}{3};\\ x,&\frac{1}{3}\leq x\leq\frac{2}{3};\\ 1,&x>\frac{2}{3}.\end{cases}$ (34)
Since the formulas for $e.cost$ and $e.lp$ in the pseudometric-weighted CC are compatible with those in the classical CC, and weights $w$ affect $\mathcal{C}$ linearly, the following lemma presented by Chawla et al. [16] remains applicable.
Lemma 5 (Lemma 5 of [16]).
Suppose $f^{+}$ is a monotonically non-decreasing piecewise convex function; $f^{-}$ is a monotonically non-decreasing piecewise concave function. Then $\mathcal{C}\geq 0$ for all possible configurations if it holds whenever:
  1. 1.
    the triangle inequality is tight for $(x_{uv},x_{vw},x_{wu})$; or
  2. 2.
    each value of $x_{e}\>(e\in\{uv,vw,wu\})$ belongs to the endpoint of the domain for some piece of the rounding function $f^{s}$, where $s$ is the sign of $e$.
Therefore, it suffices to show $6\cdot LP-ALG\geq 0$ in either of the two cases outlined above.
We introduce another useful lemma:
Lemma 6.
If $x_{uv},x_{vw},x_{wu}\in\left[\frac{1}{\alpha},1-\frac{1}{\alpha}\right]$, then $\alpha\cdot LP-ALG\geq 0$.
Proof.
$\alpha\cdot e.lp_{w}-e.cost_{w}$ $\geq\alpha(1-p_{vw}p_{wu})\min\{x,1-x\}$
$\quad-\max\{p_{vw}(1-p_{wu})+(1-p_{vw})p_{wu},(1-p_{vw})(1-p_{wu})\}$
$\geq\alpha(1-p_{vw}p_{wu})\cdot\frac{1}{\alpha}-(1-p_{vw}p_{wu})=0.$
Same inequality holds for other vertices, hence $\alpha\cdot LP-ALG\geq 0$. ∎
For notational simplicity, define $(x,y,z):=(x_{uv},x_{vw},x_{wu})$ and w.l.o.g. $x\leq y\leq z$.
A.1 Triangle inequality is tight.
In this case, $z=x+y$. As the rounding functions are defined piecewise, we can partition the entire space of possible configuration of $(x,y)$ into a finite number of regions (labeled I through VI in Figure 3) where the behaviors of $p_{uv},\,p_{vw},\,p_{wu}$ are well identified in each.
Figure 3: Configuration of regions induced from (34).
Since $f^{+}=f^{-}$, the value of $p_{e}$ is independent of the sign of $e$, therefore the formula for both $e.cost_{w}$ and $e.lp_{w}$ are independent of the sign of the edges $vw$ and $wu$. Therefore, once the region and the sign of the edge $uv$ are fixed, the formula for $e.cost_{w}$ and $e.lp_{w}$ are fully resolved as a function of $x,y,z$, as shown in Table 1.
Table 1: Formula for $e.cost$ and $e.lp$ for each configuration.
$x_{uv}=x$ $x_{vw}=y$ $x_{wu}=z=x+y$
Region Sign $e.cost_{w}$ $e.lp_{w}$ $e.cost_{u}$ $e.lp_{u}$ $e.cost_{v}$ $e.lp_{v}$
I $+$ 00 $x$ 00 $y$ 00 $z$
-- $1$ $1-x$ $1$ $1-y$ $1$ $1-z$
II $+$ $z$ $x$ $z$ $y$ 00 $z$
-- $1-z$ $1-x$ $1-z$ $1-y$ $1$ $1-z$
III $+$ $y+z-2yz$ $(1-yz)x$ $z$ $y$ $y$ $z$
-- $(1-y)(1-z)$ $(1-yz)(1-x)$ $1-z$ $1-y$ $1-y$ $1-z$
IV $+$ $1-y$ $(1-y)x$ $1$ $y$ $y$ $z$
-- 00 $(1-y)(1-x)$ 00 $1-y$ $1-y$ $1-z$
V $+$ 00 00 $1$ $y$ $1$ $z$
-- 00 00 00 $1-y$ 00 $1-z$
VI $+$ $1-y$ $(1-y)x$ $1-x$ $(1-x)y$ $x+y-2xy$ $(1-xy)z$
-- 00 $(1-y)(1-x)$ 00 $(1-x)(1-y)$ $(1-x)(1-y)$ $(1-xy)(1-z)$
By the implication of Lemma 2, it suffices to verify all possible configurations of the following procedure:
  1. 1.
    Select any region: This results in $12$ corresponding cells, with $6$ possible cases.
  2. 2.
    Select any $2$ columns: This results in $8$ corresponding cells, with $3$ possible cases.
  3. 3.
    For each column, select the sign: This results in $4$ corresponding cells, with $4$ possible cases.
  4. 4.
    The following must hold: $\alpha\cdot(e.lp_{a}+e.lp_{b})-(e.cost_{a}+e.cost_{b})\geq 0$, where $a,b$ are the vertices corresponding to the selected columns.
A.1.1 Region I, II, V
All formulas are affine; therefore, checking $\alpha\cdot(e.lp_{a}+e.lp_{b})-(e.cost_{a}+e.cost_{b})\geq 0$ at the extremal points is sufficient. Table 2 shows extremal points of the regions with the value of $6\cdot e.cost-e.lp$ for each configuration. In each region with the point, adding any two values from different columns results in a nonnegative value, validating the nonnegativity throughout the region with sign configuration.
Table 2: Value of $6\cdot e.lp-e.cost$ for each extremal point of the region.
Region $(x,y)$ Sign $6\cdot e.lp_{w}-e.cost_{w}$ $6\cdot e.lp_{u}-e.cost_{u}$ $6\cdot e.lp_{v}-e.cost_{v}$
I $(0,0)$ $+$ 00 00 00
-- $5$ $5$ $5$
$(0,\frac{1}{3})$ $+$ 00 $2$ $2$
-- $5$ $3$ $3$
$(\frac{1}{6},\frac{1}{6})$ $+$ $1$ $1$ $2$
-- $4$ $4$ $3$
II $(0,\frac{1}{3})$ $+$ $-1/3$ $5/3$ $2$
-- $16/3$ $10/3$ $3$
$(\frac{1}{6},\frac{1}{6})$ $+$ $2/3$ $2/3$ $2$
-- $13/3$ $1/3$ $3$
$(\frac{1}{3},\frac{1}{3})$ $+$ $4/3$ $4/3$ $4$
-- $11/3$ $11/3$ $1$
V $(0,\frac{2}{3})$ $+$ 00 $3$ $3$
-- 00 $2$ $2$
$(0,1)$ $+$ 00 $5$ $5$
-- 00 00 00
$(\frac{1}{3},\frac{2}{3})$ $+$ 00 $3$ $5$
-- 00 $2$ 00
A.1.2 Region III
We can also further optimize procedures 3 and 4: Compute $\alpha\cdot e.lp_{a}-e.cost_{a}$ for each region, sign, and vertex configuration. For a given region and vertex, if one sign configuration dominates another, choose the smaller one.
For example, consider $6\cdot e.lp_{w}-e.cost_{w}$ in Table 1. Since $x\leq\frac{1}{2}$, $(1-yz)(1-x)\geq(1-yz)x$. For $e.cost_{w}$, $y+z-2yz\geq(1-y)(1-z)$ since
$y+z-2yz-(1-y)(1-z)=\frac{1-(2-3y)(2-3z)}{3}\geq 0$
since $y,z\in[1/3,2/3]$ within the region. Therefore, $6\cdot e.lp_{w}-e.cost_{w}$ is smaller with the sign ‘$+$’.
Now, consider the minimum value possible for the formula above within the region. Replacing $x=z-y$,
$6\cdot e.lp_{w}-e.cost_{w}=-6yz^{2}+6y^{2}z+2yz-7y+5z.$
Fixing $y$, the formula is concave of $z$; therefore, it is minimized in $z=y$ or $z=2/3$, resulting in $2y^{2}-2y$ and $4y^{2}-\frac{25}{3}y+\frac{10}{3}$ each. Within the range $y\in[1/3,2/3]$, the minimum value of each is $-1/2$ in $y=1/2$ and $-4/9$ in $y=2/3$. Thus, $26\cdot e.lp_{w}-e.cost_{w}\geq-1/$ 2.
Finally, consider the minimum value possible for $6\cdot e.lp_{u}-e.cost_{u}$ and $6\cdot e.lp_{v}-e.cost_{v}$. Again, these are affine; hence checking on extremal points is sufficient. It turns out that $36\cdot e.lp_{u}-e.cost_{u}\geq 4/$ 3, minimized at $(y,z)=(1/3,2/3)$ with ‘$+$’ sign; $36\cdot e.lp_{v}-e.cost_{v}\geq 4/$ 3, minimized at $(y,z)=(1/3,2/3)$ with ‘--’ sign.
Therefore, any objective value is greater than or equal to $4/3-1/2=5/6\geq 0$ in the region.
A.1.3 Region IV
We can utilize the optimization method explained in A.1.2 again: Since $(1-y)x\leq(1-y)(1-x)$ and $1-y\geq 0$, the formula with the sign ‘$+$’ is smaller for $6\cdot e.lp_{w}-e.cost_{w}$, which is $(6x-1)(1-y)$. This has a minimum value of $-1/3$ in $(x,y)=(0,2/3)$.
Also, since $26z-y=6x+5y\geq 10/3>5/$ 2 in this region, the formula with the sign ‘--’ is smaller for $6\cdot e.lp_{v}-e.cost_{v}$, which is $5-6x-5y$. This has a minimum value of $-1/3$ in $(x,y)=(1/3,2/3)$.
The formula for $6\cdot e.lp_{u}-e.cost_{u}$ is $16y-$ 1 or $6(1-y)$. Each of them has a minimum value $1$ in $(x,y)=(1/3,1/3)$, $2$ in $(x,y)=(0,2/3)$, respectively. Thus, $(6\cdot e.lp_{w}-e.cost_{w})+(6\cdot e.lp_{u}-e.cost_{u})$ and $(6\cdot e.lp_{u}-e.cost_{u})+(6\cdot e.lp_{v}-e.cost_{v})$ are at least $-1/3+1=2/3\geq 0$.
Finally,
$(6\cdot e.lp_{w}-e.cost_{w})+(6\cdot e.lp_{v}-e.cost_{v})$ $\geq(6x-1)(1-y)+(5-6x-5y)$
$=-6xy-4y+4\geq-6y+4\geq 0,$
Making the approximation factor $\alpha=6$ tight at $(x,y)=(1/3,2/3)$ as well as the objective value nonnegative.
A.1.4 Region VI
As in IV, the formula with sign ‘$+$’ is smaller for $6\cdot e.lp_{w}-e.cost_{w}$, which is $(6x-1)(1-y)$. This has a minimum value of $1/3$ in $(x,y)=(1/3,2/3)$.
The formula for $6\cdot e.lp_{u}-e.cost_{u}$ is $(1-x)(6y-1)$ or $6(1-x)(1-y)$. Each of them has a minimum value $2/3$ in $(x,y)=(1/3,1/3)$, $4/3$ in $(x,y)=(1/3,2/3)$, respectively.
For $6\cdot e.lp_{v}-e.cost_{v}$, consider the difference between those of ‘$+$’ sign and ‘--’ sign.
$6(1-xy)(z-(1-z))-(x+y-2xy-(1-x)(1-y))$ $\geq 2(1-xy)-(-3xy+2x+2y-1)$
$=xy-2x-2y+3$
$=(2-x)(2-y)-1\geq\frac{11}{9}>0,$
since $z\geq 2/3$. Therefore, the formula with sign ‘--’ is smaller, which is $6(1-xy)(1-z)-(1-x)(1-y)$.
Define $p=xy$, then
$6(1-xy)(1-z)-(1-x)(1-y)=6pz-5z-7p+5,$
with region $z\in[2/3,1],\,p\in\left[\frac{z}{3}-\frac{1}{9},\left(\frac{z}{2}\right)^{2}\right]$. Fixing the value of $z$, it is a decreasing function with respect to $p$; fixing the value of $p$, it is a decreasing function with respect to $z$. Therefore, the function is minimized at $z=1,\,p=\frac{1}{4}$ with the value $-\frac{1}{4}$.
Therefore, any objective value is greater than or equal to $1/3-1/4=1/12\geq 0$ in the region.
Overall, $6\cdot LP-ALG\geq 0$ when the triangle inequality is tight.
A.2 All coordinates belong to endpoint of segments.
There are total $4$ possible LP values: 0, 1/3, 2/3, 10132310,\,1/3,\,2/3,\,10 , 1 / 3 , 2 / 3 , 1. Among all the possible $(x,y,z)$ tuples, there are only $7$ tuples whose triangle inequality is not tight. Moreover, $3$ tuples $(x,y,z)=(\frac{1}{3},\frac{1}{3},\frac{1}{3}),\,(\frac{1}{3},\frac{2}{3},\frac{2}{3}),\,(\frac{2}{3},\frac{2}{3},\frac{2}{3})$ among them only consist of value $1/3$ or $2/3$, thus $3\cdot LP-ALG\geq 0$ by Lemma 6. Therefore, we only need to verify the remaining $3$ tuples.
Since left-side and right-side limits of the value of rounding function at such point might differ, we have to consider all possible limits of $6\cdot LP-ALG$ up to $82^{3}=$ 8 directions. Table 3 shows remaining $3$ tuples and the directional limit of the value of $6\cdot e.cost-e.lp$ for each configuration. Note that since $x\leq y\leq z$, only up to $4$ out of $8$ directions are required to verification.
Table 3: Value of $6\cdot e.lp-e.cost$ for each points in consideration of Case 2.
$(x,y,z)$ Sign $6\cdot e.lp_{w}-e.cost_{w}$ $6\cdot e.lp_{u}-e.cost_{u}$ $6\cdot e.lp_{v}-e.cost_{v}$
$(\frac{2}{3}-\delta,\frac{2}{3}-\delta,1)$ $+$ $1$ $1$ $26/9$
-- $2/3$ $2/3$ $-1/9$
$(\frac{2}{3}-\delta,\frac{2}{3}+\delta,1)$ $+$ 00 $1$ $5/3$
-- 00 $2/3$ 00
$(\frac{2}{3}+\delta,\frac{2}{3}+\delta,1)$ $+$ 00 00 00
-- 00 00 00
$(\frac{1}{3}-\delta,1,1)$ $+$ 00 $5$ $5$
-- 00 00 00
$(\frac{1}{3}+\delta,1,1)$ $+$ 00 $10/3$ $10/3$
-- 00 00 00
$(\frac{2}{3}-\delta,1,1)$ $+$ 00 $5/3$ $5/3$
-- 00 00 00
$(\frac{2}{3}+\delta,1,1)$ $+$ 00 00 00
-- 00 00 00
$(1,1,1)$ $+$ 00 00 00
-- 00 00 00
As in argument on region I, II, V of Case 1, in each point with direction specified, adding any two values with different column results in a nonnegative value, validating the nonnegativity.
This completes our case analysis.
Appendix B Proof for Theorem 2
We can apply the same argument with Appendix A, with the partition by regions (Figure 4) and formula for $e.cost$ and $e.lp$ (Table 4) changed accordingly.
Figure 4: Configuration of regions induced from (11).
Table 4: Formula for $e.cost$ and $e.lp$ for each configuration.
$x_{uv}=x$ $x_{vw}=y$
Region Sign $e.cost$ $e.lp$ $e.cost$ $e.lp$
I $+$ 00 $x$ 00 $y$
-- $1$ $1-x$ $1$ $1-y$
II $+$ $\frac{5}{3}z$ $x$ $\frac{5}{3}z$ $y$
-- $1-\frac{5}{3}z$ $1-x$ $1-\frac{5}{3}z$ $1-y$
III $+$ $1$ $x$ $1$ $y$
-- 00 $1-x$ 00 $1-y$
IV $+$ $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz$ $(1-\frac{25}{9}yz)x$ $\frac{5}{3}z$ $y$
-- $(1-\frac{5}{3}y)(1-\frac{5}{3}z)$ $(1-\frac{25}{9}yz)(1-x)$ $1-\frac{5}{3}z$ $1-y$
V $+$ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ $1$ $y$
-- 00 $(1-\frac{5}{3}y)(1-x)$ 00 $1-y$
VI $+$ 00 00 $1$ $y$
-- 00 00 00 $1-y$
VII $+$ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ $1-\frac{5}{3}x$ $(1-\frac{5}{3}x)y$
-- 00 $(1-\frac{5}{3}y)(1-x)$ 00 $(1-\frac{5}{3}x)(1-y)$
$x_{wu}=z=x+y$
Region Sign $e.cost$ $e.lp$
I $+$ 00 $z$
-- $1$ $1-z$
II $+$ 00 $z$
-- $1$ $1-z$
III $+$ 00 $z$
-- $1$ $1-z$
IV $+$ $\frac{5}{3}y$ $z$
-- $1-\frac{5}{3}y$ $1-z$
V $+$ $\frac{5}{3}y$ $z$
--$1-\frac{5}{3}y$ $1-z$
VI $+$ $1$ $z$
-- 00 $1-z$
VII $+$ $\frac{5}{3}x+\frac{5}{3}y-\frac{50}{9}xy$ $(1-\frac{25}{9}xy)z$
-- $(1-\frac{5}{3}x)(1-\frac{5}{3}y)$ $(1-\frac{25}{9}xy)(1-z)$
Define $(x,y,z):=(x_{uv},x_{vw},x_{wu})$ and w.l.o.g. $x\leq y\leq z$.
B.1 Triangle inequality is tight.
B.1.1 Region I, II, III, VI
Affine formulae; refer to Table 5. Also note that the factor $10/3$ is tight for $(0,\frac{2}{5},\frac{2}{5})$ with (+,+,+)(+,+,+)( + , + , + ),
Table 5: Value of $\frac{10}{3}\cdot e.lp-e.cost$ for each extremal point of the region.
Region $(x,y)$ Sign $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$
I $(0,0)$ $+$ 00 00 00
-- $7/3$ $7/3$ $7/3$
$(0,\frac{2}{5})$ $+$ 00 $4/3$ $4/3$
-- $7/3$ $1$ $1$
$(\frac{1}{5},\frac{1}{5})$ $+$ $2/3$ $2/3$ $4/3$
-- $5/3$ $5/3$ $1$
II $(0,\frac{2}{5})$ $+$ $-2/3$ $2/3$ $4/3$
-- $3$ $5/3$ $1$
$(\frac{1}{5},\frac{1}{5})$ $+$ 00 00 $4/3$
-- $7/3$ $7/3$ $1$
$(\frac{1}{5},\frac{2}{5})$ $+$ $-1/3$ $1/3$ $2$
-- $8/3$ $2$ $1/3$
$(\frac{3}{10},\frac{3}{10})$ $+$ 00 00 $2$
-- $7/3$ $7/3$ $1/3$
III $(\frac{1}{5},\frac{2}{5})$ $+$ $-1/3$ $1/3$ $2$
-- $8/3$ $2$ $1/3$
$(\frac{3}{10},\frac{3}{10})$ $+$ 00 00 $2$
-- $7/3$ $7/3$ $1/3$
$(\frac{2}{5},\frac{2}{5})$ $+$ $1/3$ $1/3$ $8/3$
-- $2$ $2$ $-1/3$
VI $(0,\frac{3}{5})$ $+$ 00 $1$ $1$
-- 00 $4/3$ $4/3$
$(0,1)$ $+$ 00 $7/3$ $7/3$
-- 00 00 00
$(\frac{2}{5},\frac{3}{5})$ $+$ 00 $1$ $7/3$
-- 00 $4/3$ 00
B.1.2 Region IV
The argument here is analogous with A.1.2: $(1-\frac{25}{9}yz)x\leq(1-\frac{25}{9}yz)(1-x)$ as $x\leq\frac{1}{2}$; $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz\geq(1-\frac{5}{3}y)(1-\frac{5}{3}z)$ as
$\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz-\left(1-\frac{5}{3}y\right)\left(1-\frac{5}{3}z\right)=\frac{1-(2-5y)(2-5z)}{3}$
and $y,z\in\left[2/5,3/5\right]$. Therefore, $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘$+$’.
Replacing $x=z-y$,
$\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}=-\frac{250}{27}yz^{2}+\frac{250}{27}y^{2}z+\frac{50}{9}yz-5y+\frac{5}{3}z.$
Fixing $y$, this formula is concave of $z$, hence minimized in either $z=y$ or $z=3/5$, which yield $\frac{50}{9}y^{2}-\frac{10}{3}y$ and $\frac{50}{9}y^{2}-5y+1$ each. Within the range $y\in[2/5,3/5]$, the minimum value of each is $-1/9$ in $y=2/5$ and $-1/8$ in $y=9/20$. Thus, $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}\geq-1/8$.
Finally, $\frac{10}{3}\cdot e.lp_{u}-e.cost_{y}\geq 1/3$, minimized at $(y,z)=(2/5,3/5)$ with ‘$+$’ sign; $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}\geq 2/3$, minimized at $(y,z)=(2/5,2/5)$ with ‘$+$’ sign.
Therefore, any objective value is at least $1/3-1/8=5/24\geq 0$ in the region.
B.1.3 Region V
The argument here is analogous with A.1.3: $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘$+$’, which is $(\frac{10}{3}x-1)(1-\frac{5}{3}y)$. This has a minimum value of $-1/9$ in $(x,y)=(1/5,2/5)$.
The formula for $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$ is either $5/3(2x+y)$ or $7/3-5/3(2x+y)$, whose global minimum value is 00 in $(x,y)=(2/5,3/5)$ with sign ‘--’.
The formula for $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ is either $110/3\cdot y-$ 1 or $10/3(1-y)$, whose global minimum value is $1/3$ in $y=2/5$ with sign ‘$+$’. Thus, $(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w})+(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})$ and $(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})+(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v})$ are at least $-1/9+1/3=2/9\geq 0$.
Finally,
$\left(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}\right)+\left(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}\right)$
$\quad\geq\min\left\{\left(\frac{10}{3}x-1\right)\left(1-\frac{5}{3}y\right)+\frac{5}{3}(2x+y),\left(\frac{10}{3}x-1\right)\left(1-\frac{5}{3}y\right)+\frac{7}{3}-\frac{5}{3}(2x+y)\right\}$
$\quad=\min\left\{\left(\frac{10}{3}x-2\right)\left(2-\frac{5}{3}y\right)+3,\frac{50}{9}xy+\frac{4}{3}\right\}$
$\quad\geq\min\left\{1,\frac{4}{3}\right\}\geq 1\geq 0.$
B.1.4 Region VII
The argument here is analogous with A.1.4: $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘$+$’, which is $(\frac{10}{3}x-1)(1-\frac{5}{3}y)$. This has a minimum value of 00 in $(x,y)=(0.4,0.6)$.
The formula for $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ is either $(1-\frac{5}{3}x)(\frac{10}{3}y-1)$ or $\frac{10}{3}(1-\frac{5}{3}x)(1-y)$, whose global minimum value is $1/9$ in $(x,y)=(2/5,2/5),(1/2,1/2)$ with sign ‘$+$’.
For $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$,
$\frac{10}{3}\left(1-\frac{25}{9}xy\right)(2z-1)-\left(\frac{5}{3}x+\frac{5}{3}y-\frac{50}{9}xy-\left(1-\frac{5}{3}x\right)\left(1-\frac{5}{3}y\right)\right)$
$\quad\geq 2\left(1-\frac{25}{9}xy\right)-\left(-\frac{25}{3}xy+\frac{10}{3}x+\frac{10}{3}y-1\right)$
$\quad=\frac{25}{9}xy-\frac{10}{3}x-\frac{10}{3}y+3$
$\quad=\left(2-\frac{5}{3}x\right)\left(2-\frac{5}{3}y\right)-1\geq\frac{1}{3}\geq 0,$
since $z\geq\frac{4}{5}$. Therefore, the formula with sign ‘--’ is smaller, which is $\frac{10}{3}(1-\frac{25}{9}xy)(1-z)-(1-\frac{5}{3}x)(1-\frac{5}{3}y)$.
Define $p=xy$, then
$\frac{10}{3}\left(1-\frac{25}{9}xy\right)(1-z)-\left(1-\frac{5}{3}x\right)\left(1-\frac{5}{3}y\right)=\frac{250}{27}pz-\frac{5}{3}z-\frac{325}{27}p+\frac{7}{3},$
with region $z\in[4/5,1],\,p\in\left[\frac{2}{5}z-\frac{4}{25},\left(\frac{z}{2}\right)^{2}\right]$. Fixing the value of $z$,it is a decreasing function with respect to $p$. Therefore, the function is minimized when $p=\left(\frac{z}{2}\right)^{2}$, which now becomes:
$\frac{125}{54}z^{3}-\frac{325}{108}z^{2}-\frac{5}{3}z+\frac{7}{3}.$
This function decreases in the range $z\in[4/5,1]$, hence the minimum value is $-1/36$ in $z=1$. Therefore, $(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w})+(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})$ and $(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})+(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v})$ are at least $1/9-1/36=1/12\geq 0$.
Finally,
$\left(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}\right)+\left(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}\right)$
$\quad\geq\left(\frac{10}{3}x-1\right)\left(1-\frac{5}{3}y\right)+\frac{10}{3}\left(1-\frac{25}{9}xy\right)(1-z)-\left(1-\frac{5}{3}x\right)\left(1-\frac{5}{3}y\right)$
$\quad=(5x-2)\left(1-\frac{5}{3}y\right)+\frac{10}{3}\left(1-\frac{25}{9}xy\right)(1-z)$
$\quad\geq 0\cdot 0+\frac{10}{3}\cdot\frac{11}{36}\cdot 0=0,$
since $x\geq\frac{2}{5},\,y\leq\frac{3}{5},\,z\leq 1$.
B.2 All coordinates belong to endpoint of segments.
There are also $4$ tuples of $(x,y,z)$ to verify, which are fully covered in Table 6. Note that there are now only single point $2/5$ where rounding functions are not continuous.
Table 6: Value of $\frac{10}{3}\cdot e.lp-e.cost$ for each points in consideration of Case 2.
$(x,y,z)$ Sign $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$
$(\frac{3}{5},\frac{3}{5},1)$ $+$ 00 00 00
-- 00 00 00
$(\frac{2}{5}-\delta,1,1)$ $+$ 00 $7/3$ $7/3$
-- 00 00 00
$(\frac{2}{5}+\delta,1,1)$ $+$ 00 $7/9$ $7/9$
-- 00 00 00
$(\frac{3}{5},1,1)$ $+$ 00 00 00
-- 00 00 00
$(1,1,1)$ $+$ 00 00 00
-- 00 00 00
This completes our case analysis.

### v2

@@
 While classical CC has been widely studied, many real-world applications involve more nuanced relationships, either multi-class categorical interactions or varying confidence levels in edge labels.
-To address these, two natural generalizations have been proposed: Chromatic Correlation Clustering (CCC), which assigns semantic colors to edge labels, and pseudometric-weighted CC, which allows edge weights satisfying the triangle inequality.
+To address these, two natural generalizations have been proposed: Chromatic Correlation Clustering, which assigns semantic colors to edge labels, and pseudometric-weighted Correlation Clustering, which allows edge weights satisfying the triangle inequality.
 In this paper, we develop improved approximation algorithms for both settings.
@@
 Clustering is a fundamental task in unsupervised learning, where the goal is to partition a set of objects into groups based on their pairwise relationships.
-One prominent problem in this domain is Correlation Clustering (CC) bansal2004correlation , which models binary similarity/dissimilarity between items using an edge-labeled graph: similar pairs are marked with a ‘$+$’ label and dissimilar pairs with a ‘--’.
+One prominent problem in this domain is Correlation Clustering (CC) bansal2004correlation , which models binary similarity/dissimilarity between items using an edge-labeled graph: similar pairs are marked with a ‘++’ label and dissimilar pairs with a ‘-’.
 The objective is to partition the nodes to minimize disagreements—i.e., cases where the partitioning contradicts the edge labels.
@@
 1.1 Related Works
-The Correlation Clustering (CC) problem has been widely studied since its introduction ben-dor1999clustering , and it is known to be APX-hard, leading to efforts to develop approximation algorithms.
+The Correlation Clustering problem has been widely studied since its introduction ben-dor1999clustering , and it is known to be APX-hard, leading to efforts to develop approximation algorithms.
 Early work by Bansal, Blum, and Chawla introduced a constant-factor approximation algorithm bansal2004correlation .
@@
 In a more recent work cao2024clusterlp , they introduced a new approach to find a feasible solution for the cluster LP in sublinear time.
-Chromatic Correlation Clustering (CCC) is an extension of the classical Correlation Clustering (CC) problem, where edge colors represent different types of relationships.
+Chromatic Correlation Clustering is an extension of the classical Correlation Clustering problem, where edge colors represent different types of relationships.
 Bonchi et al.
-bonchi2012chromatic introduced CCC and proposed a heuristic algorithm without approximation guarantees.
+bonchi2012chromatic introduced CCC with a heuristic lacking guarantees.
 Anava et al.
-anava2015chromatic presented a 4-approximation algorithm based on LP rounding and also proposed two practical methods: Reduce and Cluster (RC), with an approximation ratio of 11, and a heuristic named Deep Cluster (DC).
+anava2015chromatic gave a 4-approximation via LP rounding, plus two practical methods: Reduce and Cluster (RC, ratio 11) and Deep Cluster (DC).
 Klodt et al.
-klodt2021color showed that the Pivot algorithm by Ailon et al.
-ailon2008aggregating , originally designed for standard CC, provides a 3-approximation guarantee for CCC.
-They also established that the RC algorithm achieves a 5-approximation for CCC.
-However, Pivot’s color-agnostic nature limits its performance in practice.
-To address this, Klodt et al.
-proposed a heuristic called Greedy Expansion (GE), which performs well empirically despite lacking theoretical guarantees.
+klodt2021color showed that Pivot ailon2008aggregating yields a 3-approximation and that RC achieves 5, but Pivot’s color-agnostic design limits performance; they instead proposed Greedy Expansion (GE), a heuristic effective in practice without guarantees.
 More recently, Xiu et al.
@@
 We further prove that this approximation factor is optimal within the standard LP relaxation framework combined advanced rounding functions.
-  • For the Chromatic Correlation Clustering (CCC) problem, we enhance the LP-based method through a new analysis that yields a $2.15$-approximation, improving upon the previous best bound of $2.5$ by Xiu et al.
+  • For the Chromatic Correlation Clustering problem, we enhance the LP-based method through a new analysis that yields a $2.15$-approximation, improving upon the previous best bound of $2.5$ by Xiu et al.
 XiuHTCH22 .
@@
 2 Preliminaries
-The correlation clustering (CC) problem takes as input a signed undirected graph $G=(V,E=E^{+}\uplus E^{-})$, where each edge $e=uv\in E$ is assigned a sign ‘$+$’ or ‘--’, described by $e\in E^{+}$ or $e\in E^{-}$.
+The correlation clustering (CC) problem takes as input a signed undirected graph $G=(V,E=E^{+}\uplus E^{-})$, where each edge $e=uv\in E$ is assigned a sign ‘++’ or ‘-’, described by $e\in E^{+}$ or $e\in E^{-}$.
 The objective is to find a partition of the nodes such that the number of disagreements—i.e., negative edges within the same cluster and positive edges between different clusters—is minimized.
@@
 Since the $x$ above satisfies the triangle inequality, we can relax the range of $x$ from $\{0,1\}$ to $[0,1]$, resulting in the following LP:
-minimize $\sum_{uv\in E^{+}}x_{uv}+\sum_{uv\in E^{-}}(1-x_{uv})$ (CC-LP)
+minimize $\sum_{uv\in E^{+}}x_{uv}+\sum_{uv\in E^{-}}(1-x_{uv})$ ($C$-LP)
 subject to $x_{uv}+x_{vw}\geq x_{wu},$ (1)
@@
 2.1 Pseudometric-weighted Correlation Clustering
-The weighted Correlation Clustering (CC) problem is a generalization of the classical CC problem in which each edge is associated with a nonnegative violation cost.
-Specifically, for each edge $uv$, a weight $w_{uv}\geq 0$ is provided, and violating the edge’s label (either ‘$+$’ or ‘--’) incurs a penalty of $w_{uv}$.
+The weighted Correlation Clustering problem is a generalization of the classical CC problem in which each edge is associated with a nonnegative violation cost.
+Specifically, for each edge $uv$, a weight $w_{uv}\geq 0$ is provided, and violating the edge’s label (either ‘++’ or ‘-’) incurs a penalty of $w_{uv}$.
 This differs from the standard setting, where all violations incur a uniform cost of $1$.
@@
 In this pseudometric-weighted setting, a constant-factor approximation is known DBLP:conf/soda/CharikarG24 .
-Following the analysis of Charikar and Gao with $L=2$ yields an approximation factor of $\overline{B}_{\text{HR}}+\frac{1}{\frac{1}{3}}\leq\frac{4}{\frac{1}{3}}+2(L-1)+\frac{1}{\frac{1}{3}}=17,$ since the second type of charge occurs at most $L-1=1$ time in the charging scheme.
+Following the analysis of Charikar and Gao with $L=2$ yields an approximation factor of $\overline{B}_{\text{HR}}+\frac{1}{\frac{1}{3}}\leq\frac{4}{\frac{1}{3}}+2(L-1)+\frac{1}{\frac{1}{3}}=17,$ since the second type of charge occurs at most L1=1L-1=1 time in the charging scheme.
 The following is a natural LP relaxation of the weighted CC problem, extending (CC-LP):
@@
 2.2 Chromatic Correlation Clustering Problem
-The Chromatic Correlation Clustering (CCC) problem is a variant of the classical CC problem in which each cluster is additionally assigned a color bonchi2012chromatic .
-The input includes a set of $L$ possible colors, as well as a special color $\gamma$ that denotes that two vertices should not be placed in the same cluster—analogous to a negative (‘--’) edge in the classical CC setting.
+The Chromatic Correlation Clustering problem is a variant of the classical CC problem in which each cluster is additionally assigned a color bonchi2012chromatic .
+The input includes a set of $L$ possible colors, as well as a special color $\gamma$ that denotes that two vertices should not be placed in the same cluster—analogous to a negative (‘-’) edge in the classical CC setting.
 When $L=1$ (i.e., a single cluster color), CCC reduces to the standard CC problem.
@@
 This view generalizes the classical CC setting, where the cluster-wise discrete metric can be regarded as a special case of bottleneck distances.
-Since CCC generalizes the standard CC problem, the integrality gap of (CCC-LP) is at least as large as that of (CC-LP), which is 2.
+Since CCC generalizes the standard CC problem, the integrality gap of (CCC-LP) is at least as large as that of ($C$-LP), which is 2.
 3 Approximation Algorithm
@@
 Set $C=\{v\}$.
-for $u\in V\backslash\{v\}$do
+for $u\in V\backslash\{v\}$, do
   Set $p_{uv}$ as following:
@@
 Algorithm 2 LP-CCC
-Input: Graph $G=(V,\,E)$, color function $\phi:E\rightarrow L\cup\{\gamma\}$, LP solution $\{x_{u}^{c}\}_{u\in V,\,c\in L}$ and $\{x_{uv}^{c}\}_{uv\in E,\,c\in L}$.
+Input: Graph G=(V,E)G=(V,\,E), color function $\phi:E\rightarrow L\cup\{\gamma\}$, LP solution $\{x_{u}^{c}\}_{u\in V,\,c\in L}$ and $\{x_{uv}^{c}\}_{uv\in E,\,c\in L}$.
 Output: Clustering $\mathcal{C}$ of $V$, Coloring function $\Phi:\mathcal{C}\rightarrow L$.
@@
 for $u\in V$ do
-     if $\exists c\in L$ s.t. $x_{u}^{c}<\frac{1}{2}$then
+  if $\exists c\in L$ s.t. $x_{u}^{c}<\frac{1}{2}$, then
    Update $S_{c}\leftarrow S_{c}\cup\{u\}$.
@@
   $G_{c}=(S_{c},\,E_{c}=E_{c}^{+}\uplus E_{c}^{-}\uplus E_{c}^{\circ})$, where $E_{c}=E\cap\binom{S_{c}}{2}$,
-     and $E^{+}\uplus E^{-}\uplus E^{\circ}$ is defined as a partition by color $c,\,\gamma,L\backslash\{c\}$ respectively.
+[…]
   Set $\mathcal{C}_{c}=\textsc{LP-Pivot}(G_{c},x^{c}|_{E_{c}})$.
@@
 Rounding functions $f^{+},\,f^{-},\,f^{\circ}:[0,1]\rightarrow[0,1]$ convert the LP value $x_{uv}$ to the non-selection probability $p_{uv}$ chawla2015near .
-The sign of the edge $uv$—either ‘$+$’, ‘--’, or $\circ$’—determines which rounding function is applied.
+The sign of the edge uvuv—either ‘++’, ‘-’, or $\circ$’—determines which rounding function is applied.
 The sign $\circ$’ indicates that the edge does not belong to $E$.
@@
 When the number of distinct pairwise distances between elements, denoted by $L$, is equal to 2, their algorithm can be viewed as a special case of our LP-Pivot framework.
-In this setting, the two distances $d_{1}$ and $d_{2}$ correspond to the ‘--’ and ‘$+$’ labels, respectively, used in our rounding procedure.
+In this setting, the two distances $d_{1}$ and $d_{2}$ correspond to the ‘-’ and ‘++’ labels, respectively, used in our rounding procedure.
 The rounding functions used are:
-[…]
+$f^{+}(x)=f^{-}(x)=\begin{cases}0,&x<\alpha;\\ \frac{\max\{x-\alpha\beta,\,0\}}{1-\alpha\beta},&\alpha\leq x\leq 1-\alpha;\\ 1,&x>1-\alpha,\end{cases}\qquad f^{\circ}(x)=\begin{cases}0,&x<\alpha\beta;\\ x,&\alpha\beta\leq x\leq 1-\alpha\beta;\\ 1,&x>1-\alpha\beta.\end{cases}$
 Here, $\alpha$ and $\beta$ are fixed algorithmic parameters. For the pseudometric-weighted CC problem, they are set to $\alpha=\frac{1}{3}$ and $\beta=0$, and $f^{\circ}$ is unused.
-With triple-based analysis, this choice yields an approximation factor of $6$, as shown in the Appendix.
+With triple-based analysis, this choice yields an approximation factor of $6$, as shown in the subsection 4.4.
 Chromatic CC: The LP-based pivoting algorithm by Xiu et al. XiuHTCH22 uses LP values directly as probabilities, corresponding to the following rounding functions:
@@
 This setting is known to achieve an approximation factor of $2.5$.
+4.4 Proof for $6$-approximation on Pseudometric-weighted CC with LP-UMVD-Pivot
+With the parameters $\alpha=\frac{1}{3}$ and $\beta=0$, the corresponding rounding functions are as follows:
+$f^{+}(x)=f^{-}(x)=\begin{cases}0,&x<\frac{1}{3};\\ x,&\frac{1}{3}\leq x\leq\frac{2}{3};\\ 1,&x>\frac{2}{3}.\end{cases}$ (14)
+Since the formulas for $e.cost$ and $e.lp$ in the pseudometric-weighted CC are compatible with those in the classical CC, and weights $w$ affect $\mathcal{C}$ linearly, the following lemma presented by Chawla et al.
+chawla2015near remains applicable.
+Lemma 2 (Lemma 5 of chawla2015near ).
+Suppose $f^{+}$ is a monotonically non-decreasing piecewise convex function; $f^{-}$ is a monotonically non-decreasing piecewise concave function.
+Then $\mathcal{C}\geq 0$ for all possible configurations if it holds whenever:
+  1. 1.
+    the triangle inequality is tight for $(x_{uv},x_{vw},x_{wu})$; or
+  2. 2.
+    each value of $x_{e}\>(e\in\{uv,vw,wu\})$ belongs to the endpoint of the domain for some piece of the rounding function $f^{s}$, where $s$ is the sign of $e$.
+Therefore, it suffices to show $6\cdot LP-ALG\geq 0$ in either of the two cases outlined above.
+We introduce another useful lemma:
+Lemma 3.
+If $x_{uv},x_{vw},x_{wu}\in\left[\frac{1}{\alpha},1-\frac{1}{\alpha}\right]$, then $\alpha\cdot LP-ALG\geq 0$.
+Proof.
+$\alpha\cdot e.lp_{w}-e.cost_{w}$ $\geq\alpha(1-p_{vw}p_{wu})\min\{x,1-x\}$
+$\quad-\max\{p_{vw}(1-p_{wu})+(1-p_{vw})p_{wu},(1-p_{vw})(1-p_{wu})\}$
+$\geq\alpha(1-p_{vw}p_{wu})\cdot\frac{1}{\alpha}-(1-p_{vw}p_{wu})=0.$
+Same inequality holds for other vertices, hence $\alpha\cdot LP-ALG\geq 0$. ∎
+For notational simplicity, define $(x,y,z):=(x_{uv},x_{vw},x_{wu})$ and w.l.o.g. $x\leq y\leq z$.
+4.4.1 Triangle inequality is tight.
+In this case, $z=x+y$.
+As the rounding functions are defined piecewise, we can partition the entire space of possible configuration of $(x,y)$ into a finite number of regions (labeled I through VI in Figure 2) where the behaviors of $p_{uv},\,p_{vw},\,p_{wu}$ are well identified in each.
+Figure 2: Configuration of regions induced from (14).
+Since $f^{+}=f^{-}$, the value of $p_{e}$ is independent of the sign of $e$, therefore the formula for both $e.cost_{w}$ and $e.lp_{w}$ are independent of the sign of the edges $vw$ and $wu$.
+Therefore, once the region and the sign of the edge $uv$ are fixed, the formula for $e.cost_{w}$ and $e.lp_{w}$ are fully resolved as a function of $x,y,z$, as shown in Table 1.
+Table 1: Formula for $e.cost$ and $e.lp$ for each configuration.
+$x_{uv}=x$ $x_{vw}=y$ $x_{wu}=z=x+y$
+Region Sign $e.cost_{w}$ $e.lp_{w}$ $e.cost_{u}$ $e.lp_{u}$ $e.cost_{v}$ $e.lp_{v}$
+I ++ 0 $x$ 0 $y$ 0 $z$
+- 11 1x1-x 11 1y1-y 11 1z1-z
+II ++ $z$ $x$ $z$ $y$ 0 $z$
+- 1z1-z 1x1-x 1z1-z 1y1-y 11 1z1-z
+III ++ y+z2yzy+z-2yz (1yz)x(1-yz)x $z$ $y$ $y$ $z$
+- (1y)(1z)(1-y)(1-z) (1yz)(1x)(1-yz)(1-x) 1z1-z 1y1-y 1y1-y 1z1-z
+IV ++ 1y1-y (1y)x(1-y)x 11 $y$ $y$ $z$
+- 0 (1y)(1x)(1-y)(1-x) 0 1y1-y 1y1-y 1z1-z
+V ++ 0 0 11 $y$ 11 $z$
+- 0 0 0 1y1-y 0 1z1-z
+VI ++ 1y1-y (1y)x(1-y)x 1x1-x (1x)y(1-x)y x+y2xyx+y-2xy (1xy)z(1-xy)z
+- 0 (1y)(1x)(1-y)(1-x) 0 (1x)(1y)(1-x)(1-y) (1x)(1y)(1-x)(1-y) (1xy)(1z)(1-xy)(1-z)
+By the implication of Lemma 4, it suffices to verify all possible configurations of the following procedure:
+  1. 1.
+    Select any region: This results in $12$ corresponding cells, with $6$ possible cases.
+  2. 2.
+    Select any $2$ columns: This results in $8$ corresponding cells, with $3$ possible cases.
+  3. 3.
+    For each column, select the sign: This results in $4$ corresponding cells, with $4$ possible cases.
+  4. 4.
+    The following must hold: $\alpha\cdot(e.lp_{a}+e.lp_{b})-(e.cost_{a}+e.cost_{b})\geq 0$, where $a,b$ are the vertices corresponding to the selected columns.
+Region I, II, V
+All formulas are affine; therefore, checking $\alpha\cdot(e.lp_{a}+e.lp_{b})-(e.cost_{a}+e.cost_{b})\geq 0$ at the extremal points is sufficient.
+Table 2 shows extremal points of the regions with the value of $6\cdot e.cost-e.lp$ for each configuration.
+In each region with the point, adding any two values from different columns results in a nonnegative value, validating the nonnegativity throughout the region with sign configuration.
+Table 2: Value of $6\cdot e.lp-e.cost$ for each extremal point of the region.
+Region $(x,y)$ Sign $6\cdot e.lp_{w}-e.cost_{w}$ $6\cdot e.lp_{u}-e.cost_{u}$ $6\cdot e.lp_{v}-e.cost_{v}$
+I $(0,0)$ ++ 0 0 0
+- 55 55 55
+$(0,\frac{1}{3})$ ++ 0 22 22
+- 55 33 33
+$(\frac{1}{6},\frac{1}{6})$ ++ 11 11 22
+- 44 44 33
+II $(0,\frac{1}{3})$ ++ 1/3-1/3 $5/3$ 22
+- $16/3$ $10/3$ 33
+$(\frac{1}{6},\frac{1}{6})$ ++ $2/3$ $2/3$ 22
+- $13/3$ $1/3$ 33
+$(\frac{1}{3},\frac{1}{3})$ ++ $4/3$ $4/3$ 44
+- $11/3$ $11/3$ 11
+V $(0,\frac{2}{3})$ ++ 0 33 33
+- 0 22 22
+$(0,1)$ ++ 0 55 55
+- 0 0 0
+$(\frac{1}{3},\frac{2}{3})$ ++ 0 33 55
+- 0 22 0
+Region III
+We can also further optimize procedures 3 and 4: Compute $\alpha\cdot e.lp_{a}-e.cost_{a}$ for each region, sign, and vertex configuration.
+For a given region and vertex, if one sign configuration dominates another, choose the smaller one.
+For example, consider $6\cdot e.lp_{w}-e.cost_{w}$ in Table 1. Since $x\leq\frac{1}{2}$, $(1-yz)(1-x)\geq(1-yz)x$. For $e.cost_{w}$, $y+z-2yz\geq(1-y)(1-z)$ since
+$y+z-2yz-(1-y)(1-z)=\frac{1-(2-3y)(2-3z)}{3}\geq 0$
+since $y,z\in[1/3,2/3]$ within the region. Therefore, $6\cdot e.lp_{w}-e.cost_{w}$ is smaller with the sign ‘++’.
+Now, consider the minimum value possible for the formula above within the region. Replacing x=zyx=z-y,
+$6\cdot e.lp_{w}-e.cost_{w}=-6yz^{2}+6y^{2}z+2yz-7y+5z.$
+Fixing $y$, the formula is concave of $z$; therefore, it is minimized in $z=y$ or $z=2/3$, resulting in $2y^{2}-2y$ and $4y^{2}-\frac{25}{3}y+\frac{10}{3}$ each.
+Within the range $y\in[1/3,2/3]$, the minimum value of each is 1/2-1/2 in $y=1/2$ and 4/9-4/9 in $y=2/3$.
+Thus, $6\cdot e.lp_{w}-e.cost_{w}\geq-1/2$.
+Finally, consider the minimum value possible for $6\cdot e.lp_{u}-e.cost_{u}$ and $6\cdot e.lp_{v}-e.cost_{v}$.
+Again, these are affine; hence checking on extremal points is sufficient.
+It turns out that $6\cdot e.lp_{u}-e.cost_{u}\geq 4/3$, minimized at $(y,z)=(1/3,2/3)$ with ‘++’ sign; $6\cdot e.lp_{v}-e.cost_{v}\geq 4/3$, minimized at $(y,z)=(1/3,2/3)$ with ‘-’ sign.
+Therefore, any objective value is greater than or equal to $4/3-1/2=5/6\geq 0$ in the region.
+Region IV
+We can utilize the optimization method explained in 4.4 again: Since $(1-y)x\leq(1-y)(1-x)$ and $1-y\geq 0$, the formula with the sign ‘++’ is smaller for $6\cdot e.lp_{w}-e.cost_{w}$, which is (6x1)(1y)(6x-1)(1-y).
+This has a minimum value of 1/3-1/3 in $(x,y)=(0,2/3)$.
+Also, since $6z-y=6x+5y\geq 10/3>5/2$ in this region, the formula with the sign ‘-’ is smaller for $6\cdot e.lp_{v}-e.cost_{v}$, which is 56x5y5-6x-5y.
+This has a minimum value of 1/3-1/3 in $(x,y)=(1/3,2/3)$.
+The formula for $6\cdot e.lp_{u}-e.cost_{u}$ is 6y16y-1 or 6(1y)6(1-y).
+Each of them has a minimum value $1$ in $(x,y)=(1/3,1/3)$, 22 in $(x,y)=(0,2/3)$, respectively.
+Thus, $(6\cdot e.lp_{w}-e.cost_{w})+(6\cdot e.lp_{u}-e.cost_{u})$ and $(6\cdot e.lp_{u}-e.cost_{u})+(6\cdot e.lp_{v}-e.cost_{v})$ are at least $1/3+1=2/3\geq 0$.
+Finally,
+$(6\cdot e.lp_{w}-e.cost_{w})+(6\cdot e.lp_{v}-e.cost_{v})$ $\geq(6x-1)(1-y)+(5-6x-5y)$
+$=-6xy-4y+4\geq-6y+4\geq 0,$
+Making the approximation factor $\alpha=6$ tight at $(x,y)=(1/3,2/3)$ as well as the objective value nonnegative.
+Region VI
+As in IV, the formula with sign ‘++’ is smaller for $6\cdot e.lp_{w}-e.cost_{w}$, which is (6x1)(1y)(6x-1)(1-y). This has a minimum value of $1/3$ in $(x,y)=(1/3,2/3)$.
+The formula for $6\cdot e.lp_{u}-e.cost_{u}$ is (1x)(6y1)(1-x)(6y-1) or 6(1x)(1y)6(1-x)(1-y). Each of them has a minimum value $2/3$ in $(x,y)=(1/3,1/3)$, $4/3$ in $(x,y)=(1/3,2/3)$, respectively.
+For $6\cdot e.lp_{v}-e.cost_{v}$, consider the difference between those of ‘++’ sign and ‘-’ sign.
+$6(1-xy)(z-(1-z))-(x+y-2xy-(1-x)(1-y))$ $\geq 2(1-xy)-(-3xy+2x+2y-1)$
+$=xy-2x-2y+3$
+$=(2-x)(2-y)-1\geq\frac{11}{9}>0,$
+since $z\geq 2/3$. Therefore, the formula with sign ‘-’ is smaller, which is 6(1xy)(1z)(1x)(1y)6(1-xy)(1-z)-(1-x)(1-y).
+Define $p=xy$, then
+6(1xy)(1z)(1x)(1y)=6pz5z7p+5,6(1-xy)(1-z)-(1-x)(1-y)=6pz-5z-7p+5,
+with region $z\in[2/3,1],\,p\in\left[\frac{z}{3}-\frac{1}{9},\left(\frac{z}{2}\right)^{2}\right]$.
+Fixing the value of $z$, it is a decreasing function with respect to $p$; fixing the value of $p$, it is a decreasing function with respect to $z$.
+Therefore, the function is minimized at $z=1,\,p=\frac{1}{4}$ with the value $\frac{1}{4}$.
+Therefore, any objective value is greater than or equal to $1/3-1/4=1/12\geq 0$ in the region.
+Overall, $6\cdot LP-ALG\geq 0$ when the triangle inequality is tight.
+4.4.2 All coordinates belong to endpoint of segments.
+There are total $4$ possible LP values: 0, 1/3, 2/3, 10,\,1/3,\,2/3,\,1.
+Among all the possible $(x,y,z)$ tuples, there are only $7$ tuples whose triangle inequality is not tight.
+Moreover, $3$ tuples $(x,y,z)=(\frac{1}{3},\frac{1}{3},\frac{1}{3}),\,(\frac{1}{3},\frac{2}{3},\frac{2}{3}),\,(\frac{2}{3},\frac{2}{3},\frac{2}{3})$ among them only consist of value $1/3$ or $2/3$, thus $3\cdot LP-ALG\geq 0$ by Lemma 3.
+Therefore, we only need to verify the remaining $3$ tuples.
+Since left-side and right-side limits of the value of rounding function at such point might differ, we have to consider all possible limits of $6\cdot LP-ALG$ up to $2^{3}=8$ directions.
+Table 3 shows remaining $3$ tuples and the directional limit of the value of $6\cdot e.cost-e.lp$ for each configuration.
+Note that since $x\leq y\leq z$, only up to $4$ out of $8$ directions are required to verification.
+Table 3: Value of $6\cdot e.lp-e.cost$ for each points in consideration of Case 2.
+$(x,y,z)$ Sign $6\cdot e.lp_{w}-e.cost_{w}$ $6\cdot e.lp_{u}-e.cost_{u}$ $6\cdot e.lp_{v}-e.cost_{v}$
+$(\frac{2}{3}-\delta,\frac{2}{3}-\delta,1)$ ++ 11 11 $26/9$
+- $2/3$ $2/3$ 1/9-1/9
+$(\frac{2}{3}-\delta,\frac{2}{3}+\delta,1)$ ++ 0 11 $5/3$
+- 0 $2/3$ 0
+$(\frac{2}{3}+\delta,\frac{2}{3}+\delta,1)$ ++ 0 0 0
+- 0 0 0
+$(\frac{1}{3}-\delta,1,1)$ ++ 0 55 55
+- 0 0 0
+$(\frac{1}{3}+\delta,1,1)$ ++ 0 $10/3$ $10/3$
+- 0 0 0
+$(\frac{2}{3}-\delta,1,1)$ ++ 0 $5/3$ $5/3$
+- 0 0 0
+$(\frac{2}{3}+\delta,1,1)$ ++ 0 0 0
+- 0 0 0
+$(1,1,1)$ ++ 0 0 0
+- 0 0 0
+As in argument on region I, II, V of Case 1, in each point with direction specified, adding any two values with different column results in a nonnegative value, validating the nonnegativity.
+This completes our case analysis.
 5 Triple-based Analysis
@@
 and
-$e.cost_{w}(u,v)$ =$=\begin{cases}p_{uw}(1-p_{vw})+(1-p_{uw})p_{vw},&uv\in E^{+};\\ (1-p_{uw})(1-p_{vw}),&uv\in E^{-},\end{cases}$
-$e.lp_{w}(u,v)$ =$=\begin{cases}(1-p_{uw}p_{vw})x_{uv},&uv\in E^{+};\\ (1-p_{uw}p_{vw})(1-x_{uv}),&uv\in E^{-}.\end{cases}$
+$e.cost_{w}(u,v)$ $=\begin{cases}p_{uw}(1-p_{vw})+(1-p_{uw})p_{vw},&uv\in E^{+};\\ (1-p_{uw})(1-p_{vw}),&uv\in E^{-},\end{cases}$
+$e.lp_{w}(u,v)$ $=\begin{cases}(1-p_{uw}p_{vw})x_{uv},&uv\in E^{+};\\ (1-p_{uw}p_{vw})(1-x_{uv}),&uv\in E^{-}.\end{cases}$
 In the weighted CC setting, edge weights further influence the value of $\mathcal{C}$:
@@
 Therefore, the function $\alpha LP(uvw)-ALG(uvw)$ is linear w.r.p. $(w_{uv},w_{vw},w_{wu})$.
-Since the set of $(w_{uv},w_{vw},w_{wu})$ that satisfies the triangle inequality forms a convex cone generated by $(1,1,0),\,(1,0,1),\,(0,1,1)$, the function value is nonnegative for all such $(w_{uv},w_{vw},w_{wu})$ if and only if the value is nonnegative for $(w_{uv},w_{vw},w_{wu})\in\{(1,1,0),\,(1,0,1),\,(0,1,1)\}$.
+Since the set of $(w_{uv},w_{vw},w_{wu})$ that satisfies the triangle inequality forms a convex cone generated by (1,1,0),(1,0,1),(0,1,1)(1,1,0),\,(1,0,1),\,(0,1,1), the function value is nonnegative for all such $(w_{uv},w_{vw},w_{wu})$ if and only if the value is nonnegative for $(w_{uv},w_{vw},w_{wu})\in\{(1,1,0),\,(1,0,1),\,(0,1,1)\}$.
 ∎
-This lemma implies that the algorithm achieves an approximation factor of $\alpha$ if all of the following inequalities are satisfied for every possible configuration on the triangle $uvw$:
+This lemma implies that the algorithm achieves an approximation factor of $\alpha$ if all of the following inequalities are satisfied for every possible configuration on the triangle uvwuvw:
 $e.cost_{w}(uv)+e.cost_{u}(vw)$ $\leq\alpha\cdot(e.lp_{w}(uv)+e.lp_{u}(vw)),$
@@
 We obtain a lower bound on the approximation factor of LP-Pivot by verifying the feasibility of rounding functions that satisfy the above inequalities.
-To this end, we analyze several configurations of LP values and edge signs on triangle $uvw$.
+To this end, we analyze several configurations of LP values and edge signs on triangle uvwuvw.
 In Theorems 1 and 3, the notation ‘$(a,b,c)$ with $(s_{1},s_{2},s_{3})$’ denotes $(x_{uv},x_{vw},x_{wu})=(a,b,c)$, where each edge sign is given by $s_{1}$, $s_{2}$, and $s_{3}$, respectively.
@@
 Suppose that LP-Pivot achieves an expected $\alpha$-approximation.
-We begin by considering the configuration $(x,1-x,1)$ with signs (+,+,)(+,+,-)( + , + , - ) or (+,,)(+,-,-)( + , - , - ) and weights $(1,0,1)$. Then:
+We begin by considering the configuration (x,1x,1)(x,1-x,1) with signs (+,+,)(+,+,-) or (+,,)(+,-,-) and weights $(1,0,1)$. Then:
 $\alpha(1-f^{+}(1-x))x$ $\geq(2-f^{+}(x))(1-f^{+}(1-x)),$ (15)
@@
 $f^{+}(x)\leq 2-\alpha x\quad\implies\quad f^{+}(1-x)=f^{-}(1-x)=1.$ (17)
-Next, consider the configuration $(0,x,x)$ with signs (+,+,+)(+,+,+)( + , + , + ) and weights $(1,1,0)$:
+Next, consider the configuration $(0,x,x)$ with signs $(+,+,+)$ and weights $(1,1,0)$:
 $\alpha x\geq 3f^{+}(x)-2f^{+}(x)^{2},$ (18)
@@
 $f^{+}(x)\leq\alpha x\leq 2-\alpha x,\quad f^{+}(1-x)=1.$
-Now consider the configuration $(x,x,2x)$ with signs (+,+,+)(+,+,+)( + , + , + ) and weights $(1,1,0)$:
+Now consider the configuration $(x,x,2x)$ with signs $(+,+,+)$ and weights $(1,1,0)$:
 $\alpha x(1-f^{+}(x)f^{+}(2x))\geq f^{+}(x)+f^{+}(2x)-2f^{+}(x)f^{+}(2x).$ (19)
@@
 These formulations are central to the analysis. Since $\alpha\cdot LP-ALG$ is always at least the expression obtained from the LP lower bound, we can prove that this bound is nonnegative.
-As in Section 5.1, the algorithm achieves an $\alpha$-approximation if the following inequality holds for all triangles $uvw$:
+As in Section 5.1, the algorithm achieves an $\alpha$-approximation if the following inequality holds for all triangles uvwuvw:
 $e.cost_{w}(uv)+e.cost_{u}(vw)+e.cost_{v}(wu)\leq\alpha\cdot\left(e.lp_{w}(uv)+e.lp_{u}(vw)+e.lp_{v}(wu)\right).$
@@
 Proof.
-Suppose LP-CCC is an expected $\alpha$-approximation algorithm, where $42\leq\alpha<$ 4.
-Plugging $(\frac{1}{2},\frac{1}{2},1)$ with (,,)($\circ$,$\circ$,-)( ∘ , ∘ , - ), we obtain:
+Suppose LP-CCC is an expected $\alpha$-approximation algorithm, where $2\leq\alpha<4$.
+Plugging $(\frac{1}{2},\frac{1}{2},1)$ with $(\circ,\circ,-)$, we obtain:
 $\alpha\left(1-f^{\circ}\left(\frac{1}{2}\right)\right)\geq 2\left(1-f^{\circ}\left(\frac{1}{2}\right)\right)+\left(1-f^{\circ}\left(\frac{1}{2}\right)\right)^{2},$ (24)
@@
 $f^{\circ}\left(\frac{1}{2}\right)\geq 3-\alpha.$ (25)
-Next, plugging $\left(\frac{1}{4},\frac{1}{4},\frac{1}{2}\right)$ with (+,+,)(+,+$\circ$)( + , + , ∘ ) yields:
+Next, plugging $\left(\frac{1}{4},\frac{1}{4},\frac{1}{2}\right)$ with $(+,+,\circ)$ yields:
 $\frac{1-f^{+}\left(\frac{1}{4}\right)^{2}+2\left(f^{+}\left(\frac{1}{4}\right)+f^{\circ}\left(\frac{1}{2}\right)-2f^{+}\left(\frac{1}{4}\right)f^{\circ}\left(\frac{1}{2}\right)\right)}{\frac{3}{4}\left(1-f^{+}\left(\frac{1}{4}\right)^{2}\right)+\frac{1}{2}\left(1-f^{+}\left(\frac{1}{4}\right)f^{\circ}\left(\frac{1}{2}\right)\right)}\leq\alpha,$ (26)
@@
 $\left(2+\left(\frac{\alpha}{2}-4\right)f^{+}\left(\frac{1}{4}\right)\right)f^{\circ}\left(\frac{1}{2}\right)\leq\frac{\alpha}{4}\left(5-3f^{+}\left(\frac{1}{4}\right)^{2}\right)+\left(f^{+}\left(\frac{1}{4}\right)-1\right)^{2}-2.$ (27)
-Using the known bound from chawla2015near , plugging $(0,x,x)$ with (+,+,+)(+,+,+)( + , + , + ) yields:
+Using the known bound from chawla2015near , plugging $(0,x,x)$ with $(+,+,+)$ yields:
 $f^{+}(x)\leq 1-\sqrt{1-\alpha x},$ (28)
@@
 This is the upper bound for $f^{\circ}\left(\frac{1}{2}\right)$ when $\alpha=4(1-t^{2})$.
-On the other hand, from (24), the lower bound is $13-\alpha=3-4(1-t^{2})=4t^{2}-$ 1.
+On the other hand, from (25), the lower bound is $3-\alpha=3-4(1-t^{2})=4t^{2}-1$.
 Therefore, the inequality
-$62(t^{2}-t+1)(4t^{2}-1)\leq 3t^{3}-6t^{2}-4t+$ 6 (33)
+$2(t^{2}-t+1)(4t^{2}-1)\leq 3t^{3}-6t^{2}-4t+6$ (34)
 must hold.
@@
 $h(t):=2(t^{2}-t+1)(4t^{2}-1)-(3t^{3}-6t^{2}-4t+6)=8t^{4}-11t^{3}+12t^{2}+6t-8.$
-Since $h^{\prime\prime}(t)=96t^{2}-66t+24>0$ and $h^{\prime}(0)=6>0$, $h$ is convex and increasing for $t\geq 0$.
+Since $h^{\prime\prime}(t)=96t^{2}-66t+24>0$ and $^{\prime}(0)=6>0, hh$ is convex and increasing for $t\geq 0$.
 Finally, since $h\left(\sqrt{1-\frac{2.11}{4}}\right)>0$, we conclude that $h(t)>0$ when $t\geq\sqrt{1-\frac{2.11}{4}}$, implying that the lower bound for $f^{\circ}\left(\frac{1}{2}\right)$ exceeds the upper bound when $\alpha\leq 2.11$.
@@
 Therefore, Lemma 2 does not apply, and we must instead verify the inequality $\alpha\cdot LP-ALG\geq 0$ directly.
-For sign configurations that do not involve a $\circ$’ (i.e., all edges have $+$ or -- signs), the inequality $\alpha\cdot LP-ALG\geq 0$ with $\alpha=2.06-\varepsilon$ is already established in chawla2015near .
+For sign configurations that do not involve a $\circ$’ (i.e., all edges have ++ or - signs), the inequality $\alpha\cdot LP-ALG\geq 0$ with $\alpha=2.06-\varepsilon$ is already established in chawla2015near .
 Thus, it suffices to check the remaining $6$ configurations involving at least one $\circ$’ sign.
 As before, define $(x,y,z):=(x_{uv},x_{vw},x_{wu})$ and $(p_{x},p_{y},p_{z}):=(p_{uv},p_{vw},p_{wu})$.
-5.3.1 (,,)($\circ$,$\circ$,$\circ$)( ∘ , ∘ , ∘ ) Triangles
+5.3.1 $(\circ,\circ,\circ)$ Triangles
 $2\cdot e.lp_{w}(u,v)\geq 2(1-p_{uw}p_{vw})\cdot\frac{1}{2}=(1-p_{uw}p_{vw})=e.cost_{w}(u,v).$
 Since the same inequality holds for all permutations of the endpoints, we conclude:
-$2\cdot LP-ALG\geq 0$
-5.3.2 (,,+)($\circ$,$\circ$,+)( ∘ , ∘ , + ) Triangles
+$2\cdot LP-ALG\geq 0.$
+5.3.2 $(\circ,\circ,+)$ Triangles
 If $z\geq\frac{1}{2}$,
@@
 $=(1-2z)(1-p_{x})(1-p_{y})+(1-p_{x})(1-p_{y})\geq 0.$
-5.3.3 (,,)($\circ$,$\circ$,-)( ∘ , ∘ , - ) Triangles
+5.3.3 $(\circ,\circ,-)$ Triangles
 If $z\leq 1-\frac{1}{\alpha}$, then as in Section 5.3.2, we have $\alpha\cdot LP-ALG\geq 0$.
-Otherwise, since $21-z\leq 1/\alpha<1/$ 2,
+Otherwise, since $1-z\leq 1/\alpha<1/2$,
 $\alpha\cdot LP-ALG$
@@
     $=(\alpha(1-x)-1)(2-p_{x}-p_{y})-(1-p_{x})(1-p_{y}).$
-    Note that the function is affine with respect to $p_{y}$; therefore, the function is minimized when $p_{y}$ reaches the boundary, i.e., $y=1$ or $y=1-x$.
+    Note that the function is affine with respect to $p_{y}$; therefore, the function is minimized when $p_{y}$ reaches the boundary, i.e., $y=1$ or y=1xy=1-x.
     If $y=1$, then $p_{y}=1$, hence
@@
     $\geq\left(\frac{\alpha}{2}-1\right)(1-p_{x})\geq 0.$
-    On the other hand, if $y=1-x$, then $p_{y}=\frac{3}{10}y+\frac{7}{10}=1-\frac{3}{10}x$ as $y>1/2$, hence
+    On the other hand, if y=1xy=1-x, then $p_{y}=\frac{3}{10}y+\frac{7}{10}=1-\frac{3}{10}x$ as $y>1/2$, hence
     $(\alpha(1-x)-1)(2-p_{x}-p_{y})-(1-p_{x})(1-p_{y})$
@@
     since $\frac{dp_{y}}{dy}\leq 1.7<\alpha$. Hence, the function is minimized when $y$ is maximized.
-    Applying $y=1-x$, since $y=1-x>1/2$, $p_{y}=\frac{3}{10}y+\frac{7}{10}=1-\frac{3}{10}x$ and thus
+    Applying y=1xy=1-x, since y=1x>1/2y=1-x>1/2, $p_{y}=\frac{3}{10}y+\frac{7}{10}=1-\frac{3}{10}x$ and thus
     $0.3\cdot(\alpha(1-x)-1)+\alpha(1-p_{x}p_{y})(1-x-y)-(1-p_{x})(1-p_{y})$
@@
     since $x<1/2$.
-5.3.4 (,+,+)($\circ$,+,+)( ∘ , + , + ) Triangles
+5.3.4 $(\circ,+,+)$ Triangles
 For Sections 5.3.4 and 5.3.5, we utilize the following lemma, which serves as a weaker version of Lemma 2:
@@
 $\quad+\alpha(y+z)-(1+p_{y}+p_{z}-p_{y}p_{z})$
-The first term is a monotonically decreasing function of $x$; the second term is a monotonically decreasing function of $p_{x}$ since $-\alpha(yp_{z}+p_{y}z)-2+2(p_{y}+p_{z})\leq 0$, as shown in the following Lemma:
+The first term is a monotonically decreasing function of $x$; the second term is a monotonically decreasing function of $p_{x}$ since $\alpha(yp_{z}+p_{y}z)-2+2(p_{y}+p_{z})\leq 0$, as shown in the following Lemma:
 Lemma 6.
@@
 $\alpha(xf^{+}(y)+f^{+}(x)y)+2-2(f^{+}(x)+f^{+}(y))\geq 0$
-in $[0,1]^{2}$.
+in $0,1]^{2}$.
 Proof.
@@
     is concave of $x$ since $\alpha\cdot 0.5095-2<0$ and $f^{+}$ is convex in the range.
-Hence, this is minimized at either $x=0$ or $x=0.5095$, where the former case results in 00 and the latter case results in $2\alpha\cdot 0.5095-2\geq 0$.
+Hence, this is minimized at either $x=0$ or $x=0.5095$, where the former case results in 0 and the latter case results in $2\alpha\cdot 0.5095-2\geq 0$.
   2. 2.
@@
     $=2.225\left(\left(\frac{6}{89}\right)^{2}-\left(p_{y}-\frac{83}{89}\right)\left(p_{z}-\frac{83}{89}\right)\right).$
-    Since $z\geq 1/2$, $p_{z}>0.94>\frac{83}{89}$, thus the function is decreasing with respect to $p_{y}$. The minimum value of the function is therefore 00 when $p_{y}=p_{z}=1$.
+    Since $z\geq 1/2$, $p_{z}>0.94>\frac{83}{89}$, thus the function is decreasing with respect to $p_{y}$. The minimum value of the function is therefore 0 when $p_{y}=p_{z}=1$.
   2. 2.
     $y+z<1$
-    We use the similar argument with of those conducted by Chawla et al.chawla2015near on the (+,+,)(+,+,-)( + , + , - ) case. For simplicity, let $a=0.19,\,b=0.5095$.
+    We use the similar argument with of those conducted by Chawla et al.chawla2015near on the (+,+,)(+,+,-) case. For simplicity, let a=0.19,b=0.5095a=0.19,\,b=0.5095.
     If $y<a$, $p_{y}=0$, hence
@@
     We now show that the function is minimized when $y=z$ as well as $x=y+z$.
-Consider the parametrization $(x,y,z)=(x,t-c,t+c)$, where $t\in[a,1/2],\,c\in[0,\min\{t-a,b-t\}],$ and $x=2t$ since $\alpha\cdot LP-ALG$ is a monotonically increasing function of $x$.
+Consider the parametrization (x,y,z)=(x,tc,t+c)(x,y,z)=(x,t-c,t+c), where $t\in[a,1/2],\,c\in[0,\min\{t-a,b-t\}],$ and $x=2t$ since $\alpha\cdot LP-ALG$ is a monotonically increasing function of $x$.
 Then
@@
     $\quad+\alpha(1-p_{y})c.$
-    Fixing the value of $t$, the sum of first $2$ terms is a negative even quartic function of $c$, where the negativity is obtained from $-\alpha(1-t)+1\leq-\alpha(1-b)+1<0$.
+    Fixing the value of $t$, the sum of first $2$ terms is a negative even quartic function of $c$, where the negativity is obtained from $\alpha(1-t)+1\leq-\alpha(1-b)+1<0$.
 Thus, we can write the function as
@@
 Since our proposed $f^{\circ}$ doesn’t intersect the region, $\alpha\cdot LP-ALG\geq 0$ if $y=z$, and thus $\alpha\cdot LP-ALG\geq 0$ whenever $a\leq y\leq z\leq b$.
-5.3.5 (,+,)($\circ$,+,-)( ∘ , + , - ) Triangles
+5.3.5 $(\circ,+,-)$ Triangles
 We would prove that the lower bound $\mathcal{C}^{\prime}$ of $\mathcal{C}$, which is defined as
@@
     The first term is a monotonically decreasing function of $p_{x}$ since
-    $-\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z=(\alpha(p_{y}-y)+2)z-(\alpha+1)p_{y}$
+    $\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z=(\alpha(p_{y}-y)+2)z-(\alpha+1)p_{y}$
     is an increasing function of $z$ since $p_{y}\geq y$ in the range,
-    $-\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z\leq-\alpha y-p_{y}+2\leq-\frac{\alpha}{2}-f^{+}\left(\frac{1}{2}\right)+2\approx-0.016<0.$
+    $\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z\leq-\alpha y-p_{y}+2\leq-\frac{\alpha}{2}-f^{+}\left(\frac{1}{2}\right)+2\approx-0.016<0.$
     Hence, the function is minimized when $x$ is maximized. Setting $x=1$,
@@
     Since $y\geq 1/2$ and $p_{y}\leq 1$,
-    $-\alpha+1-\alpha y+\left(\frac{\alpha}{2}+1\right)p_{y}\leq-\alpha+2<0.$
+    $\alpha+1-\alpha y+\left(\frac{\alpha}{2}+1\right)p_{y}\leq-\alpha+2<0.$
     Hence, the function is minimized when $z$ is maximized. Setting $z=1$,
@@
     $\mathcal{C}^{\prime}$ $\geq\alpha\cdot\left((1-p_{y})(1-y)+(1-p_{x})y\right)-\left(3-2p_{x}-2p_{y}+p_{x}p_{y}\right).$
-    As mentioned above, it is sufficient to check on $x=1$ and $x=1-y$.
+    As mentioned above, it is sufficient to check on $x=1$ and x=1yx=1-y.
     1. (a)
@@
     2. (b)
-      $x=1-y$
-      Since $x=1-y>1/2$, $p_{x}=\frac{3}{10}x+\frac{7}{10}=1-\frac{3}{10}y$. Therefore,
+      x=1yx=1-y
+      Since x=1y>1/2x=1-y>1/2, $p_{x}=\frac{3}{10}x+\frac{7}{10}=1-\frac{3}{10}y$. Therefore,
       $\mathcal{C}^{\prime}$ $=\alpha\cdot\left((1-p_{y})(1-y)+\frac{3}{10}y^{2}\right)-\left(1+\frac{3}{10}y\right)(2-p_{y})+1.$
@@
       $=\frac{1}{200}(129y^{2}-550y+230)>0.65\geq 0.$
-      Otherwise, i.e. $0.50950.19\leq y<1/2<$ 0.5095,
+      Otherwise, i.e. $0.19\leq y<1/2<0.5095$,
       $\mathcal{C}^{\prime}\approx 0.74331+2.39737y-19.7409y^{2}+24.0007y^{3},$
@@
     Differentiating by $z$,
-    $1\displaystyle 2\alpha p_{y}z-\alpha(2+p_{y}-p_{x}p_{y}+p_{x}y)+2p_{x}+p_{y}-$ 1
+    $2\alpha p_{y}z-\alpha(2+p_{y}-p_{x}p_{y}+p_{x}y)+2p_{x}+p_{y}-1$
     $\leq\alpha p_{y}-\alpha(2+p_{y}-p_{x}p_{y}+p_{x}y)+2p_{x}+p_{y}-1$
     $=(-2\alpha+2p_{x}+p_{y})-\alpha((1-p_{x})p_{y}+p_{x}y)-1\leq 0$
-    since $z<1/2$ and $-2\alpha+2p_{x}+p_{y}\leq-2\alpha+3=-1.3<0$. Therefore, the function is minimized when $z$ is maximized.
+    since $z<1/2$ and $2\alpha+2p_{x}+p_{y}\leq-2\alpha+3=-1.3<0$. Therefore, the function is minimized when $z$ is maximized.
     Plugging $z=y$, this now reduce to Case 3 with $z=y$ due to the continuity of $\mathcal{C}^{\prime}$.
-5.3.6 (,,)($\circ$,-,-)( ∘ , - , - ) Triangles
+5.3.6 $(\circ,-,-)$ Triangles
 As in 5.3.1, $2e.lp_{w}(u,v)\geq e.cost_{w}(u,v)$. For other two terms,
@@
 6 Conclusion
-In this work, we studied two important variants of correlation clustering: pseudometric-weighted correlation clustering and chromatic correlation clustering (CCC).
+In this work, we studied two important variants of correlation clustering: pseudometric-weighted correlation clustering and chromatic correlation clustering .
 For both problems, we developed and analyzed specialized rounding functions that are essential for achieving improved approximation guarantees via the LP-Pivot algorithm.
 For the pseudometric-weighted setting, we proposed a piecewise-linear rounding function tailored for the setting that achieves a $10/3$-approximation, and showed that no alternative function within our analytical framework can yield a better factor.
-For the chromatic correlation clustering (CCC) variant, we designed a distinct rounding function that respects the constraints imposed by color restrictions and achieves an approximation factor of $2.15$.
+For the chromatic correlation clustering variant, we designed a distinct rounding function that respects the constraints imposed by color restrictions and achieves an approximation factor of $2.15$.
 The function is constructed using a piecewise-linear form and leverages a careful analysis of triple costs.
 Overall, our work highlights the importance of designing principled and variant-specific rounding strategies to extend LP-based techniques to structured clustering problems, yielding strong theoretical guarantees.
-Appendix A Proof for $6$-approximation on Pseudometric-weighted CC with LP-UMVD-Pivot
-With the parameters $\alpha=\frac{1}{3}$ and $\beta=0$, the corresponding rounding functions are as follows:
-$f^{+}(x)=f^{-}(x)=\begin{cases}0,&x<\frac{1}{3};\\ x,&\frac{1}{3}\leq x\leq\frac{2}{3};\\ 1,&x>\frac{2}{3}.\end{cases}$ (34)
-Since the formulas for $e.cost$ and $e.lp$ in the pseudometric-weighted CC are compatible with those in the classical CC, and weights $w$ affect $\mathcal{C}$ linearly, the following lemma presented by Chawla et al.
-[16] remains applicable.
-Lemma 5 (Lemma 5 of [16]).
-Suppose $f^{+}$ is a monotonically non-decreasing piecewise convex function; $f^{-}$ is a monotonically non-decreasing piecewise concave function.
-Then $\mathcal{C}\geq 0$ for all possible configurations if it holds whenever:
-  1. 1.
-    the triangle inequality is tight for $(x_{uv},x_{vw},x_{wu})$; or
-  2. 2.
-    each value of $x_{e}\>(e\in\{uv,vw,wu\})$ belongs to the endpoint of the domain for some piece of the rounding function $f^{s}$, where $s$ is the sign of $e$.
-Therefore, it suffices to show $6\cdot LP-ALG\geq 0$ in either of the two cases outlined above.
-We introduce another useful lemma:
-Lemma 6.
-If $x_{uv},x_{vw},x_{wu}\in\left[\frac{1}{\alpha},1-\frac{1}{\alpha}\right]$, then $\alpha\cdot LP-ALG\geq 0$.
-Proof.
-$\alpha\cdot e.lp_{w}-e.cost_{w}$ $\geq\alpha(1-p_{vw}p_{wu})\min\{x,1-x\}$
-$\quad-\max\{p_{vw}(1-p_{wu})+(1-p_{vw})p_{wu},(1-p_{vw})(1-p_{wu})\}$
-$\geq\alpha(1-p_{vw}p_{wu})\cdot\frac{1}{\alpha}-(1-p_{vw}p_{wu})=0.$
-Same inequality holds for other vertices, hence $\alpha\cdot LP-ALG\geq 0$. ∎
-For notational simplicity, define $(x,y,z):=(x_{uv},x_{vw},x_{wu})$ and w.l.o.g. $x\leq y\leq z$.
-A.1 Triangle inequality is tight.
-In this case, $z=x+y$.
-As the rounding functions are defined piecewise, we can partition the entire space of possible configuration of $(x,y)$ into a finite number of regions (labeled I through VI in Figure 3) where the behaviors of $p_{uv},\,p_{vw},\,p_{wu}$ are well identified in each.
-Figure 3: Configuration of regions induced from (34).
-Since $f^{+}=f^{-}$, the value of $p_{e}$ is independent of the sign of $e$, therefore the formula for both $e.cost_{w}$ and $e.lp_{w}$ are independent of the sign of the edges $vw$ and $wu$.
-Therefore, once the region and the sign of the edge $uv$ are fixed, the formula for $e.cost_{w}$ and $e.lp_{w}$ are fully resolved as a function of $x,y,z$, as shown in Table 1.
-Table 1: Formula for $e.cost$ and $e.lp$ for each configuration.
-$x_{uv}=x$ $x_{vw}=y$ $x_{wu}=z=x+y$
-Region Sign $e.cost_{w}$ $e.lp_{w}$ $e.cost_{u}$ $e.lp_{u}$ $e.cost_{v}$ $e.lp_{v}$
-I $+$ 00 $x$ 00 $y$ 00 $z$
--- $1$ $1-x$ $1$ $1-y$ $1$ $1-z$
-II $+$ $z$ $x$ $z$ $y$ 00 $z$
--- $1-z$ $1-x$ $1-z$ $1-y$ $1$ $1-z$
-III $+$ $y+z-2yz$ $(1-yz)x$ $z$ $y$ $y$ $z$
--- $(1-y)(1-z)$ $(1-yz)(1-x)$ $1-z$ $1-y$ $1-y$ $1-z$
-IV $+$ $1-y$ $(1-y)x$ $1$ $y$ $y$ $z$
--- 00 $(1-y)(1-x)$ 00 $1-y$ $1-y$ $1-z$
-V $+$ 00 00 $1$ $y$ $1$ $z$
--- 00 00 00 $1-y$ 00 $1-z$
-VI $+$ $1-y$ $(1-y)x$ $1-x$ $(1-x)y$ $x+y-2xy$ $(1-xy)z$
--- 00 $(1-y)(1-x)$ 00 $(1-x)(1-y)$ $(1-x)(1-y)$ $(1-xy)(1-z)$
-By the implication of Lemma 2, it suffices to verify all possible configurations of the following procedure:
-  1. 1.
-    Select any region: This results in $12$ corresponding cells, with $6$ possible cases.
-  2. 2.
-    Select any $2$ columns: This results in $8$ corresponding cells, with $3$ possible cases.
-  3. 3.
-    For each column, select the sign: This results in $4$ corresponding cells, with $4$ possible cases.
-  4. 4.
-    The following must hold: $\alpha\cdot(e.lp_{a}+e.lp_{b})-(e.cost_{a}+e.cost_{b})\geq 0$, where $a,b$ are the vertices corresponding to the selected columns.
-A.1.1 Region I, II, V
-All formulas are affine; therefore, checking $\alpha\cdot(e.lp_{a}+e.lp_{b})-(e.cost_{a}+e.cost_{b})\geq 0$ at the extremal points is sufficient.
-Table 2 shows extremal points of the regions with the value of $6\cdot e.cost-e.lp$ for each configuration.
-In each region with the point, adding any two values from different columns results in a nonnegative value, validating the nonnegativity throughout the region with sign configuration.
-Table 2: Value of $6\cdot e.lp-e.cost$ for each extremal point of the region.
-Region $(x,y)$ Sign $6\cdot e.lp_{w}-e.cost_{w}$ $6\cdot e.lp_{u}-e.cost_{u}$ $6\cdot e.lp_{v}-e.cost_{v}$
-I $(0,0)$ $+$ 00 00 00
--- $5$ $5$ $5$
-$(0,\frac{1}{3})$ $+$ 00 $2$ $2$
--- $5$ $3$ $3$
-$(\frac{1}{6},\frac{1}{6})$ $+$ $1$ $1$ $2$
--- $4$ $4$ $3$
-II $(0,\frac{1}{3})$ $+$ $-1/3$ $5/3$ $2$
--- $16/3$ $10/3$ $3$
-$(\frac{1}{6},\frac{1}{6})$ $+$ $2/3$ $2/3$ $2$
--- $13/3$ $1/3$ $3$
-$(\frac{1}{3},\frac{1}{3})$ $+$ $4/3$ $4/3$ $4$
--- $11/3$ $11/3$ $1$
-V $(0,\frac{2}{3})$ $+$ 00 $3$ $3$
--- 00 $2$ $2$
-$(0,1)$ $+$ 00 $5$ $5$
--- 00 00 00
-$(\frac{1}{3},\frac{2}{3})$ $+$ 00 $3$ $5$
--- 00 $2$ 00
-A.1.2 Region III
-We can also further optimize procedures 3 and 4: Compute $\alpha\cdot e.lp_{a}-e.cost_{a}$ for each region, sign, and vertex configuration.
-For a given region and vertex, if one sign configuration dominates another, choose the smaller one.
-For example, consider $6\cdot e.lp_{w}-e.cost_{w}$ in Table 1. Since $x\leq\frac{1}{2}$, $(1-yz)(1-x)\geq(1-yz)x$. For $e.cost_{w}$, $y+z-2yz\geq(1-y)(1-z)$ since
-$y+z-2yz-(1-y)(1-z)=\frac{1-(2-3y)(2-3z)}{3}\geq 0$
-since $y,z\in[1/3,2/3]$ within the region. Therefore, $6\cdot e.lp_{w}-e.cost_{w}$ is smaller with the sign ‘$+$’.
-Now, consider the minimum value possible for the formula above within the region. Replacing $x=z-y$,
-$6\cdot e.lp_{w}-e.cost_{w}=-6yz^{2}+6y^{2}z+2yz-7y+5z.$
-Fixing $y$, the formula is concave of $z$; therefore, it is minimized in $z=y$ or $z=2/3$, resulting in $2y^{2}-2y$ and $4y^{2}-\frac{25}{3}y+\frac{10}{3}$ each.
-Within the range $y\in[1/3,2/3]$, the minimum value of each is $-1/2$ in $y=1/2$ and $-4/9$ in $y=2/3$.
-Thus, $26\cdot e.lp_{w}-e.cost_{w}\geq-1/$ 2.
-Finally, consider the minimum value possible for $6\cdot e.lp_{u}-e.cost_{u}$ and $6\cdot e.lp_{v}-e.cost_{v}$.
-Again, these are affine; hence checking on extremal points is sufficient.
-It turns out that $36\cdot e.lp_{u}-e.cost_{u}\geq 4/$ 3, minimized at $(y,z)=(1/3,2/3)$ with ‘$+$’ sign; $36\cdot e.lp_{v}-e.cost_{v}\geq 4/$ 3, minimized at $(y,z)=(1/3,2/3)$ with ‘--’ sign.
-Therefore, any objective value is greater than or equal to $4/3-1/2=5/6\geq 0$ in the region.
-A.1.3 Region IV
-We can utilize the optimization method explained in A.1.2 again: Since $(1-y)x\leq(1-y)(1-x)$ and $1-y\geq 0$, the formula with the sign ‘$+$’ is smaller for $6\cdot e.lp_{w}-e.cost_{w}$, which is $(6x-1)(1-y)$.
-This has a minimum value of $-1/3$ in $(x,y)=(0,2/3)$.
-Also, since $26z-y=6x+5y\geq 10/3>5/$ 2 in this region, the formula with the sign ‘--’ is smaller for $6\cdot e.lp_{v}-e.cost_{v}$, which is $5-6x-5y$.
-This has a minimum value of $-1/3$ in $(x,y)=(1/3,2/3)$.
-The formula for $6\cdot e.lp_{u}-e.cost_{u}$ is $16y-$ 1 or $6(1-y)$.
-Each of them has a minimum value $1$ in $(x,y)=(1/3,1/3)$, $2$ in $(x,y)=(0,2/3)$, respectively.
-Thus, $(6\cdot e.lp_{w}-e.cost_{w})+(6\cdot e.lp_{u}-e.cost_{u})$ and $(6\cdot e.lp_{u}-e.cost_{u})+(6\cdot e.lp_{v}-e.cost_{v})$ are at least $-1/3+1=2/3\geq 0$.
-Finally,
-$(6\cdot e.lp_{w}-e.cost_{w})+(6\cdot e.lp_{v}-e.cost_{v})$ $\geq(6x-1)(1-y)+(5-6x-5y)$
-$=-6xy-4y+4\geq-6y+4\geq 0,$
-Making the approximation factor $\alpha=6$ tight at $(x,y)=(1/3,2/3)$ as well as the objective value nonnegative.
-A.1.4 Region VI
-As in IV, the formula with sign ‘$+$’ is smaller for $6\cdot e.lp_{w}-e.cost_{w}$, which is $(6x-1)(1-y)$. This has a minimum value of $1/3$ in $(x,y)=(1/3,2/3)$.
-The formula for $6\cdot e.lp_{u}-e.cost_{u}$ is $(1-x)(6y-1)$ or $6(1-x)(1-y)$. Each of them has a minimum value $2/3$ in $(x,y)=(1/3,1/3)$, $4/3$ in $(x,y)=(1/3,2/3)$, respectively.
-For $6\cdot e.lp_{v}-e.cost_{v}$, consider the difference between those of ‘$+$’ sign and ‘--’ sign.
-$6(1-xy)(z-(1-z))-(x+y-2xy-(1-x)(1-y))$ $\geq 2(1-xy)-(-3xy+2x+2y-1)$
-$=xy-2x-2y+3$
-$=(2-x)(2-y)-1\geq\frac{11}{9}>0,$
-since $z\geq 2/3$. Therefore, the formula with sign ‘--’ is smaller, which is $6(1-xy)(1-z)-(1-x)(1-y)$.
-Define $p=xy$, then
-$6(1-xy)(1-z)-(1-x)(1-y)=6pz-5z-7p+5,$
-with region $z\in[2/3,1],\,p\in\left[\frac{z}{3}-\frac{1}{9},\left(\frac{z}{2}\right)^{2}\right]$.
-Fixing the value of $z$, it is a decreasing function with respect to $p$; fixing the value of $p$, it is a decreasing function with respect to $z$.
-Therefore, the function is minimized at $z=1,\,p=\frac{1}{4}$ with the value $-\frac{1}{4}$.
-Therefore, any objective value is greater than or equal to $1/3-1/4=1/12\geq 0$ in the region.
-Overall, $6\cdot LP-ALG\geq 0$ when the triangle inequality is tight.
-A.2 All coordinates belong to endpoint of segments.
-There are total $4$ possible LP values: 0, 1/3, 2/3, 10132310,\,1/3,\,2/3,\,10 , 1 / 3 , 2 / 3 , 1.
-Among all the possible $(x,y,z)$ tuples, there are only $7$ tuples whose triangle inequality is not tight.
-Moreover, $3$ tuples $(x,y,z)=(\frac{1}{3},\frac{1}{3},\frac{1}{3}),\,(\frac{1}{3},\frac{2}{3},\frac{2}{3}),\,(\frac{2}{3},\frac{2}{3},\frac{2}{3})$ among them only consist of value $1/3$ or $2/3$, thus $3\cdot LP-ALG\geq 0$ by Lemma 6.
-Therefore, we only need to verify the remaining $3$ tuples.
-Since left-side and right-side limits of the value of rounding function at such point might differ, we have to consider all possible limits of $6\cdot LP-ALG$ up to $82^{3}=$ 8 directions.
-Table 3 shows remaining $3$ tuples and the directional limit of the value of $6\cdot e.cost-e.lp$ for each configuration.
-Note that since $x\leq y\leq z$, only up to $4$ out of $8$ directions are required to verification.
-Table 3: Value of $6\cdot e.lp-e.cost$ for each points in consideration of Case 2.
-$(x,y,z)$ Sign $6\cdot e.lp_{w}-e.cost_{w}$ $6\cdot e.lp_{u}-e.cost_{u}$ $6\cdot e.lp_{v}-e.cost_{v}$
-$(\frac{2}{3}-\delta,\frac{2}{3}-\delta,1)$ $+$ $1$ $1$ $26/9$
--- $2/3$ $2/3$ $-1/9$
-$(\frac{2}{3}-\delta,\frac{2}{3}+\delta,1)$ $+$ 00 $1$ $5/3$
--- 00 $2/3$ 00
-$(\frac{2}{3}+\delta,\frac{2}{3}+\delta,1)$ $+$ 00 00 00
--- 00 00 00
-$(\frac{1}{3}-\delta,1,1)$ $+$ 00 $5$ $5$
--- 00 00 00
-$(\frac{1}{3}+\delta,1,1)$ $+$ 00 $10/3$ $10/3$
--- 00 00 00
-$(\frac{2}{3}-\delta,1,1)$ $+$ 00 $5/3$ $5/3$
--- 00 00 00
-$(\frac{2}{3}+\delta,1,1)$ $+$ 00 00 00
--- 00 00 00
-$(1,1,1)$ $+$ 00 00 00
--- 00 00 00
-As in argument on region I, II, V of Case 1, in each point with direction specified, adding any two values with different column results in a nonnegative value, validating the nonnegativity.
-This completes our case analysis.
-Appendix B Proof for Theorem 2
-We can apply the same argument with Appendix A, with the partition by regions (Figure 4) and formula for $e.cost$ and $e.lp$ (Table 4) changed accordingly.
+Appendix A Proof for Theorem 2
+We can apply the same argument with Appendix 4.4, with the partition by regions (Figure 4) and formula for $e.cost$ and $e.lp$ (Table 4) changed accordingly.
 Figure 4: Configuration of regions induced from (11).
@@
 Region Sign $e.cost$ $e.lp$ $e.cost$ $e.lp$
-I $+$ 00 $x$ 00 $y$
--- $1$ $1-x$ $1$ $1-y$
-II $+$ $\frac{5}{3}z$ $x$ $\frac{5}{3}z$ $y$
--- $1-\frac{5}{3}z$ $1-x$ $1-\frac{5}{3}z$ $1-y$
-III $+$ $1$ $x$ $1$ $y$
--- 00 $1-x$ 00 $1-y$
-IV $+$ $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz$ $(1-\frac{25}{9}yz)x$ $\frac{5}{3}z$ $y$
--- $(1-\frac{5}{3}y)(1-\frac{5}{3}z)$ $(1-\frac{25}{9}yz)(1-x)$ $1-\frac{5}{3}z$ $1-y$
-V $+$ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ $1$ $y$
--- 00 $(1-\frac{5}{3}y)(1-x)$ 00 $1-y$
-VI $+$ 00 00 $1$ $y$
--- 00 00 00 $1-y$
-VII $+$ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ $1-\frac{5}{3}x$ $(1-\frac{5}{3}x)y$
--- 00 $(1-\frac{5}{3}y)(1-x)$ 00 $(1-\frac{5}{3}x)(1-y)$
+I ++ 0 $x$ 0 $y$
+- 11 1x1-x 11 1y1-y
+II ++ $\frac{5}{3}z$ $x$ $\frac{5}{3}z$ $y$
+- $1-\frac{5}{3}z$ 1x1-x $1-\frac{5}{3}z$ 1y1-y
+III ++ 11 $x$ 11 $y$
+- 0 1x1-x 0 1y1-y
+IV ++ $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz$ $(1-\frac{25}{9}yz)x$ $\frac{5}{3}z$ $y$
+- $(1-\frac{5}{3}y)(1-\frac{5}{3}z)$ $(1-\frac{25}{9}yz)(1-x)$ $1-\frac{5}{3}z$ 1y1-y
+V ++ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ 11 $y$
+- 0 $(1-\frac{5}{3}y)(1-x)$ 0 1y1-y
+VI ++ 0 0 11 $y$
+- 0 0 0 1y1-y
+VII ++ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ $1-\frac{5}{3}x$ $(1-\frac{5}{3}x)y$
+- 0 $(1-\frac{5}{3}y)(1-x)$ 0 $(1-\frac{5}{3}x)(1-y)$
 $x_{wu}=z=x+y$
 Region Sign $e.cost$ $e.lp$
-I $+$ 00 $z$
--- $1$ $1-z$
-II $+$ 00 $z$
--- $1$ $1-z$
-III $+$ 00 $z$
--- $1$ $1-z$
-IV $+$ $\frac{5}{3}y$ $z$
--- $1-\frac{5}{3}y$ $1-z$
-V $+$ $\frac{5}{3}y$ $z$
---$1-\frac{5}{3}y$ $1-z$
-VI $+$ $1$ $z$
--- 00 $1-z$
-VII $+$ $\frac{5}{3}x+\frac{5}{3}y-\frac{50}{9}xy$ $(1-\frac{25}{9}xy)z$
+I ++ 0 $z$
+- 11 1z1-z
+II ++ 0 $z$
+- 11 1z1-z
+III ++ 0 $z$
+- 11 1z1-z
+IV ++ $\frac{5}{3}y$ $z$
+- $1-\frac{5}{3}y$ 1z1-z
+V ++ $\frac{5}{3}y$ $z$
+-$1-\frac{5}{3}y$ 1z1-z
+VI ++ 11 $z$
+- 0 1z1-z
+VII ++ $\frac{5}{3}x+\frac{5}{3}y-\frac{50}{9}xy$ $(1-\frac{25}{9}xy)z$
 - $(1-\frac{5}{3}x)(1-\frac{5}{3}y)$ $(1-\frac{25}{9}xy)(1-z)$
 Define $(x,y,z):=(x_{uv},x_{vw},x_{wu})$ and w.l.o.g. $x\leq y\leq z$.
-B.1 Triangle inequality is tight.
-B.1.1 Region I, II, III, VI
-Affine formulae; refer to Table 5. Also note that the factor $10/3$ is tight for $(0,\frac{2}{5},\frac{2}{5})$ with (+,+,+)(+,+,+)( + , + , + ),
+A.1 Triangle inequality is tight.
+A.1.1 Region I, II, III, VI
+Affine formulae; refer to Table 5. Also note that the factor $10/3$ is tight for $(0,\frac{2}{5},\frac{2}{5})$ with $(+,+,+)$,
 Table 5: Value of $\frac{10}{3}\cdot e.lp-e.cost$ for each extremal point of the region.
 Region $(x,y)$ Sign $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$
-I $(0,0)$ $+$ 00 00 00
+I $(0,0)$ ++ 0 0 0
 - $7/3$ $7/3$ $7/3$
-$(0,\frac{2}{5})$ $+$ 00 $4/3$ $4/3$
--- $7/3$ $1$ $1$
-$(\frac{1}{5},\frac{1}{5})$ $+$ $2/3$ $2/3$ $4/3$
--- $5/3$ $5/3$ $1$
-II $(0,\frac{2}{5})$ $+$ $-2/3$ $2/3$ $4/3$
--- $3$ $5/3$ $1$
-$(\frac{1}{5},\frac{1}{5})$ $+$ 00 00 $4/3$
--- $7/3$ $7/3$ $1$
-$(\frac{1}{5},\frac{2}{5})$ $+$ $-1/3$ $1/3$ $2$
--- $8/3$ $2$ $1/3$
-$(\frac{3}{10},\frac{3}{10})$ $+$ 00 00 $2$
+$(0,\frac{2}{5})$ ++ 0 $4/3$ $4/3$
+- $7/3$ 11 11
+$(\frac{1}{5},\frac{1}{5})$ ++ $2/3$ $2/3$ $4/3$
+- $5/3$ $5/3$ 11
+II $(0,\frac{2}{5})$ ++ 2/3-2/3 $2/3$ $4/3$
+- 33 $5/3$ 11
+$(\frac{1}{5},\frac{1}{5})$ ++ 0 0 $4/3$
+- $7/3$ $7/3$ 11
+$(\frac{1}{5},\frac{2}{5})$ ++ 1/3-1/3 $1/3$ 22
+- $8/3$ 22 $1/3$
+$(\frac{3}{10},\frac{3}{10})$ ++ 0 0 22
 - $7/3$ $7/3$ $1/3$
-III $(\frac{1}{5},\frac{2}{5})$ $+$ $-1/3$ $1/3$ $2$
--- $8/3$ $2$ $1/3$
-$(\frac{3}{10},\frac{3}{10})$ $+$ 00 00 $2$
+III $(\frac{1}{5},\frac{2}{5})$ ++ 1/3-1/3 $1/3$ 22
+- $8/3$ 22 $1/3$
+$(\frac{3}{10},\frac{3}{10})$ ++ 0 0 22
 - $7/3$ $7/3$ $1/3$
-$(\frac{2}{5},\frac{2}{5})$ $+$ $1/3$ $1/3$ $8/3$
--- $2$ $2$ $-1/3$
-VI $(0,\frac{3}{5})$ $+$ 00 $1$ $1$
--- 00 $4/3$ $4/3$
-$(0,1)$ $+$ 00 $7/3$ $7/3$
--- 00 00 00
-$(\frac{2}{5},\frac{3}{5})$ $+$ 00 $1$ $7/3$
--- 00 $4/3$ 00
-B.1.2 Region IV
-The argument here is analogous with A.1.2: $(1-\frac{25}{9}yz)x\leq(1-\frac{25}{9}yz)(1-x)$ as $x\leq\frac{1}{2}$; $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz\geq(1-\frac{5}{3}y)(1-\frac{5}{3}z)$ as
+$(\frac{2}{5},\frac{2}{5})$ ++ $1/3$ $1/3$ $8/3$
+- 22 22 1/3-1/3
+VI $(0,\frac{3}{5})$ ++ 0 11 11
+- 0 $4/3$ $4/3$
+$(0,1)$ ++ 0 $7/3$ $7/3$
+- 0 0 0
+$(\frac{2}{5},\frac{3}{5})$ ++ 0 11 $7/3$
+- 0 $4/3$ 0
+A.1.2 Region IV
+The argument here is analogous with 4.4: $(1-\frac{25}{9}yz)x\leq(1-\frac{25}{9}yz)(1-x)$ as $x\leq\frac{1}{2}$; $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz\geq(1-\frac{5}{3}y)(1-\frac{5}{3}z)$ as
 $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz-\left(1-\frac{5}{3}y\right)\left(1-\frac{5}{3}z\right)=\frac{1-(2-5y)(2-5z)}{3}$
-and $y,z\in\left[2/5,3/5\right]$. Therefore, $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘$+$’.
-Replacing $x=z-y$,
+and $y,z\in\left[2/5,3/5\right]$. Therefore, $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘++’.
+Replacing x=zyx=z-y,
 $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}=-\frac{250}{27}yz^{2}+\frac{250}{27}y^{2}z+\frac{50}{9}yz-5y+\frac{5}{3}z.$
 Fixing $y$, this formula is concave of $z$, hence minimized in either $z=y$ or $z=3/5$, which yield $\frac{50}{9}y^{2}-\frac{10}{3}y$ and $\frac{50}{9}y^{2}-5y+1$ each.
-Within the range $y\in[2/5,3/5]$, the minimum value of each is $-1/9$ in $y=2/5$ and $-1/8$ in $y=9/20$.
+Within the range $y\in[2/5,3/5]$, the minimum value of each is 1/9-1/9 in $y=2/5$ and 1/8-1/8 in $y=9/20$.
 Thus, $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}\geq-1/8$.
-Finally, $\frac{10}{3}\cdot e.lp_{u}-e.cost_{y}\geq 1/3$, minimized at $(y,z)=(2/5,3/5)$ with ‘$+$’ sign; $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}\geq 2/3$, minimized at $(y,z)=(2/5,2/5)$ with ‘$+$’ sign.
+Finally, $\frac{10}{3}\cdot e.lp_{u}-e.cost_{y}\geq 1/3$, minimized at $(y,z)=(2/5,3/5)$ with ‘++’ sign; $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}\geq 2/3$, minimized at $(y,z)=(2/5,2/5)$ with ‘++’ sign.
 Therefore, any objective value is at least $1/3-1/8=5/24\geq 0$ in the region.
-B.1.3 Region V
-The argument here is analogous with A.1.3: $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘$+$’, which is $(\frac{10}{3}x-1)(1-\frac{5}{3}y)$.
-This has a minimum value of $-1/9$ in $(x,y)=(1/5,2/5)$.
-The formula for $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$ is either $5/3(2x+y)$ or $7/3-5/3(2x+y)$, whose global minimum value is 00 in $(x,y)=(2/5,3/5)$ with sign ‘--’.
-The formula for $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ is either $110/3\cdot y-$ 1 or $10/3(1-y)$, whose global minimum value is $1/3$ in $y=2/5$ with sign ‘$+$’.
-Thus, $(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w})+(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})$ and $(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})+(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v})$ are at least $-1/9+1/3=2/9\geq 0$.
+A.1.3 Region V
+The argument here is analogous with 4.4: $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘++’, which is $(\frac{10}{3}x-1)(1-\frac{5}{3}y)$.
+This has a minimum value of 1/9-1/9 in $(x,y)=(1/5,2/5)$.
+The formula for $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$ is either $5/3(2x+y)$ or 7/35/3(2x+y)7/3-5/3(2x+y), whose global minimum value is 0 in $(x,y)=(2/5,3/5)$ with sign ‘-’.
+The formula for $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ is either $10/3\cdot y-1$ or 10/3(1y)10/3(1-y), whose global minimum value is $1/3$ in $y=2/5$ with sign ‘++’.
+Thus, $(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w})+(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})$ and $(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})+(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v})$ are at least $1/9+1/3=2/9\geq 0$.
 Finally,
@@
 $\quad\geq\min\left\{1,\frac{4}{3}\right\}\geq 1\geq 0.$
-B.1.4 Region VII
-The argument here is analogous with A.1.4: $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘$+$’, which is $(\frac{10}{3}x-1)(1-\frac{5}{3}y)$.
-This has a minimum value of 00 in $(x,y)=(0.4,0.6)$.
-The formula for $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ is either $(1-\frac{5}{3}x)(\frac{10}{3}y-1)$ or $\frac{10}{3}(1-\frac{5}{3}x)(1-y)$, whose global minimum value is $1/9$ in $(x,y)=(2/5,2/5),(1/2,1/2)$ with sign ‘$+$’.
+A.1.4 Region VII
+The argument here is analogous with 4.4: $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘++’, which is $(\frac{10}{3}x-1)(1-\frac{5}{3}y)$.
+This has a minimum value of 0 in $(x,y)=(0.4,0.6)$.
+The formula for $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ is either $(1-\frac{5}{3}x)(\frac{10}{3}y-1)$ or $\frac{10}{3}(1-\frac{5}{3}x)(1-y)$, whose global minimum value is $1/9$ in $(x,y)=(2/5,2/5),(1/2,1/2)$ with sign ‘++’.
 For $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$,
@@
 $\quad=\left(2-\frac{5}{3}x\right)\left(2-\frac{5}{3}y\right)-1\geq\frac{1}{3}\geq 0,$
-since $z\geq\frac{4}{5}$. Therefore, the formula with sign ‘--’ is smaller, which is $\frac{10}{3}(1-\frac{25}{9}xy)(1-z)-(1-\frac{5}{3}x)(1-\frac{5}{3}y)$.
+since $z\geq\frac{4}{5}$. Therefore, the formula with sign ‘-’ is smaller, which is $\frac{10}{3}(1-\frac{25}{9}xy)(1-z)-(1-\frac{5}{3}x)(1-\frac{5}{3}y)$.
 Define $p=xy$, then
@@
 $\frac{125}{54}z^{3}-\frac{325}{108}z^{2}-\frac{5}{3}z+\frac{7}{3}.$
-This function decreases in the range $z\in[4/5,1]$, hence the minimum value is $-1/36$ in $z=1$.
+This function decreases in the range $z\in[4/5,1]$, hence the minimum value is 1/36-1/36 in $z=1$.
 Therefore, $(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w})+(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})$ and $(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})+(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v})$ are at least $1/9-1/36=1/12\geq 0$.
@@
 since $x\geq\frac{2}{5},\,y\leq\frac{3}{5},\,z\leq 1$.
-B.2 All coordinates belong to endpoint of segments.
+A.2 All coordinates belong to endpoint of segments.
 There are also $4$ tuples of $(x,y,z)$ to verify, which are fully covered in Table 6. Note that there are now only single point $2/5$ where rounding functions are not continuous.
@@
 $(x,y,z)$ Sign $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$
-$(\frac{3}{5},\frac{3}{5},1)$ $+$ 00 00 00
--- 00 00 00
-$(\frac{2}{5}-\delta,1,1)$ $+$ 00 $7/3$ $7/3$
--- 00 00 00
-$(\frac{2}{5}+\delta,1,1)$ $+$ 00 $7/9$ $7/9$
--- 00 00 00
-$(\frac{3}{5},1,1)$ $+$ 00 00 00
--- 00 00 00
-$(1,1,1)$ $+$ 00 00 00
--- 00 00 00
+$(\frac{3}{5},\frac{3}{5},1)$ ++ 0 0 0
+- 0 0 0
+$(\frac{2}{5}-\delta,1,1)$ ++ 0 $7/3$ $7/3$
+- 0 0 0
+$(\frac{2}{5}+\delta,1,1)$ ++ 0 $7/9$ $7/9$
+- 0 0 0
+$(\frac{3}{5},1,1)$ ++ 0 0 0
+- 0 0 0
+$(1,1,1)$ ++ 0 0 0
+- 0 0 0
 This completes our case analysis.